### v1

# TrialPanorama: Database and Benchmark for Systematic Review and Design of Clinical Trials

Abstract
Developing artificial intelligence (AI) for vertical domains requires a solid data foundation for both training and evaluation. In this work, we introduce TrialPanorama, a large-scale, structured database comprising 1,657,476 clinical trial records aggregated from 15 global sources. The database captures key aspects of trial design and execution, including trial setups, interventions, conditions, biomarkers, and outcomes, and links them to standard biomedical ontologies such as DrugBank and MedDRA. This structured and ontology-grounded design enables TrialPanorama to serve as a unified, extensible resource for a wide range of clinical trial tasks, including trial planning, design, and summarization. To demonstrate its utility, we derive a suite of benchmark tasks directly from the TrialPanorama database. The benchmark spans eight tasks across two categories: three for systematic review (study search, study screening, and evidence summarization) and five for trial design (arm design, eligibility criteria, endpoint selection, sample size estimation, and trial completion assessment). The experiments using five state-of-the-art large language models (LLMs) show that while general-purpose LLMs exhibit some zero-shot capability, their performance is still inadequate for high-stakes clinical trial workflows. We release TrialPanorama database and the benchmark to facilitate further research on AI for clinical trials.
1 Introduction
Developing new drugs is a time-consuming and costly endeavor, often requiring up to a decade and exceeding \$2 billion in expenses [1]. A substantial portion of this burden stems from the human study phase, during which clinical trials are conducted to assess safety and efficacy across phases I through IV. In addition, clinicians, researchers, and trial designers routinely search and review clinical trial records and results. These efforts support critical objectives such as designing new trials, evaluating treatments, and informing clinical decision-making in evidence-based medicine [2, 3].
Clinical trial records are dispersed across global registries, including ClinicalTrials.gov [4], the International Clinical Trials Registry Platform (ICTRP) [5], the EU Clinical Trials Register [6], and others [7, 8]. Additionally, many trials are reported in the literature, particularly on PubMed. Prior research has focused on curating subsets of this data for specific objectives, such as clinical trial results [9, 10], drug and condition annotations [11, 12], trial-publication linkages [13], and structured outcomes [14]. A unified view that integrates these diverse aspects of clinical trial data can substantially improve evidence synthesis in systematic reviews and enable more informed, data-driven approaches to designing future clinical trials [15, 16, 17]. In this work, we build TrialPanorama, a clinical trial database that integrates multiple data sources and standardizes trial-level design and result attributes, aiming to provide a great resource for clinical trial design and review as well as developing AI models for these tasks.
Artificial intelligence (AI), particularly with the emergence of large language models (LLMs), has garnered significant attention in clinical trial applications. Recent advances have explored tasks including predicting clinical trial outcomes [18, 19, 20], matching patients to eligible trials [21, 22], extracting cohort real-world data [23, 24], writing clinical trial documents [25, 26], and accelerating systematic review of the literature of clinical studies [27, 28]. However, these developments are often limited by the lack of standardized, comprehensive clinical trial data suitable for AI development. TrialPanorama bridges this gap by (1) offering curated benchmark data to support the development and evaluation of AI methods for trial applications, and (2) providing source-grounded evidence to enhance AI for trial applications through retrieval-augmented generation (RAG) [29].
Table 1: Comparison of TrialPanorama and prior clinical trial datasets and benchmarks. Columns: Size = dataset size (# trials), Src = data source (CTG = ClinicalTrials.gov), Des = structured trial design data (e.g., arms, eligibility), Res = structured trial results (e.g., adverse events), Out = trial outcomes, Rel = trial-study/publication/review relations, Ont = linked to medical ontologies (e.g., DrugBank, MedDRA), Bench-D = trial design benchmark tasks, Bench-SLR = systematic literature review benchmark tasks.
Dataset / Benchmark Size Src Des Res Out Rel Ont Bench-D Bench-SLR
TrialPanorama (Ours) 1.6M Multi(15 sources)
TOP [19] 10K CTG
CTKG [11] 8K CTG
CTGResults [9] 56K CTG
CTO [14] 125K CTG
TrialBench [30] 400K CTG
CARE [10] 700 pubs PubMed
PubMedKG [31] 480K PubMed,CTG
TrialReviewBench [28] 100 reviews PubMed,CTG
LEADSInstruct [27] 21K reviews PubMed,CTG
In this work, we introduce TrialPanorama, the largest-to-date structured clinical trial database, integrating over two million records from 16 global sources, including clinical trial registries, biomedical literature, and systematic reviews. As illustrated in Figure 1a, the database captures essential elements of clinical trials, such as study designs, interventions, conditions, outcomes, and biomarkers, and standardizes them through mappings to biomedical ontologies like DrugBank and MedDRA.
Building on this data foundation, we further demonstrate how TrialPanorama enables tasks that support AI development for trial design and systematic review. These tasks span both upstream evidence synthesis (e.g., study search, screening, summarization) and downstream trial planning (e.g., arm design, eligibility criteria, endpoint selection, sample size estimation, and trial completion prediction), as illustrated in Figures 1b and  1c. Compared to existing trial datasets and benchmarks (Table 1), TrialPanorama offers broader coverage, richer structure, and stronger alignment with real-world AI use cases in clinical research. We release the database and benchmark to facilitate future research in this space.
Figure 1: Overview of the TrialPanorama dataset and benchmark suite. (a) Schema of the structured clinical trial database, showing linked tables for trial protocols, interventions, outcomes, and metadata. (b) Summary of the systematic review benchmark tasks. (c) Summary of the clinical trial design benchmark tasks.
2 TrialPanorama: Database
As illustrated in Figure 1a, the TrialPanorama database comprises ten main tables: studies, conditions, drugs, endpoints, relations, biomarkers, outcomes, adverse events, results, and relations. These tables are organized into four conceptual clusters: trial attributes, trial protocols, trial results, and study-level links. The trial metadata cluster captures core metadata such as study title, brief summary, sponsor, start year, and recruitment status. The trial protocol cluster describes the setup and design of the study, including tested drugs, targeted conditions, and patient group allocations. The trial results cluster encompasses reported findings, including the trial outcomes, adverse events, and efficacy results. The study-level link cluster encodes relationships across studies, such as mappings between registry records and corresponding publications, or reviews aggregating multiple studies on a common clinical topic. In the following subsections, we describe the content of each table in detail. Data acquisition and processing procedures are provided in Appendix A.
2.1 Trial metadata
The studies table serves as the central registry in TrialPanorama, containing 1,657,476 clinical trial records from 15 diverse sources, including ClinicalTrials.gov, PubMed, CHICTR, etc. The distribution of the sources is illustrated in Figure 2a. Each row, uniquely indexed by a study_id, captures metadata such as trial type (INTERVENTIONAL, OBSERVATIONAL), sponsor (e.g., university, hospital, pharmaceutical), and temporal attributes including start_year, enabling longitudinal analysis. The table also includes regulatory fields like phase (e.g., PHASE1PHASE4) and recruitment status (e.g., RECRUITING, COMPLETED, SUSPENDED), providing a standardized overview of the trial’s attributes.
2.2 Trial protocols
The trial protocol cluster in TrialPanorama captures structured information describing the design and setup of clinical studies, including target populations, interventions, outcomes, and biological markers. This cluster comprises five key tables: drugs, conditions, biomarkers, dispositions, and endpoints.
Figure 2: Overview of the database statistics. a, Distribution of clinical trials across major data sources. b, Frequency distribution of the top 100 drugs mentioned in the database, with the most common drug classes annotated. c, Frequency distribution of the top 100 medical conditions, annotated by condition categories.
Drugs.
The drugs table records the therapeutic interventions evaluated across trials, comprising 1,498,985 rows and 15 columns. Each entry is linked to a study_id, with over 616,000 unique studies represented. The distribution of top frequent drugs is shown in Figure 2b. The drug_name field captures 404,175 unique names, including both generic labels (e.g., “Placebo”) and brand/composite entries. To enable structured pharmacological analysis, drugs are linked to DrugBank identifiers (drugbank_id) and normalized RxNorm terms (rx_normalized_name). Mechanism of action (MoA) information is encoded via the drug_moa_id field, which maps each drug to its biological targets, such as enzymes, receptors, or transporters, and their interaction types (e.g., inhibitor, agonist). The definitions and details of each MoA are stored in a separate drug_moa table, where drug_moa_id serves as the key linking back to the drugs table.
Conditions.
The conditions table catalogs the disease or physiological focus of each study, containing 10,133,060 entries from over 1,074,458 studies. Figure 2c shows the most frequent conditions. The condition_name field includes 5,236 unique MeSH-derived terms representing medical conditions, while condition_mesh_id ensures mapping to standardized biomedical ontologies. These annotations facilitate the retrieval and grouping of studies by clinical focus (e.g., cardiovascular or neurological diseases).
Biomarkers.
The biomarkers table records therapeutic biomolecular and clinical markers used to stratify, monitor, or predict treatment response. It contains 477,717 rows across 339,894 unique studies. Each biomarker is annotated by its biomarker_type (e.g., predictive, surrogate endpoint, safety), biomarker_name, and associated gene markers (biomarker_genes). This table supports downstream applications such as biomarker-driven subgroup analysis or treatment personalization.
Dispositions.
The dispositions table describes how participants are grouped and assigned to interventions within each study. With over 4.5 million entries covering 1.47 million unique studies, it specifies the intervention_type (e.g., DRUG, DEVICE, BEHAVIORAL), group_type (e.g., EXPERIMENTAL, COMPARATOR, PLACEBO), and intervention_name. This table is critical for reconstructing arm-level trial structure and modeling allocation strategies across diverse study designs.
Endpoints.
The endpoints table defines the primary clinical outcomes of interest across 2,621,226 rows and 1.38 million trials. Each entry includes a primary_endpoint label (e.g., Overall Survival, Quality of Life), categorized into high-level primary_endpoint_domain (e.g., mortality, physiological) and fine-grained primary_endpoint_subdomain (e.g., cardiovascular outcomes, emotional well-being).
2.3 Trial results
The trial results cluster captures the observed clinical outcomes reported at the end of each study, including both efficacy and safety measures. It comprises three main tables: outcomes, results, and adverse_events, each offering distinct perspectives on the impact of tested interventions.
Outcomes.
The outcomes table contains 532,525 entries, each uniquely mapped to a study. It records overall trial status (overall_status), such as completed, recruiting, or terminated, and provides a high-level assessment of outcome success through the outcome_type field (e.g., completed, positive outcome/primary endpoint(s) met, unknown, or terminated, other/unclear). These annotations facilitate downstream analyses such as success rate estimation and status tracking across therapeutic areas and time.
Results.
The results table comprises over 5.7 million entries covering more than 1.1 million studies. It records arm-level trial outcomes, including structured descriptions of study populations (population), tested interventions (interventions), observed clinical outcomes (outcomes), and treatment group type (group_type). The granularity of eligibility criteria and baseline characteristics captured in the population field supports fine-grained subgroup analysis, while the group_type distinction (e.g., EXPERIMENTAL, COMPARATOR) enables comparative outcome interpretation across arms.
Adverse Events.
The adverse_events table documents reported safety events across 6,145,012 rows and over 530,000 unique studies. It captures adverse_event_name (e.g., Headache, Nausea, Diarrhoea) along with corresponding MedDRA-coded identifiers (meddra_id), enabling standardized aggregation and harmonization of safety profiles across studies. These data support pharmacovigilance and risk-benefit assessments for individual therapies or therapeutic classes.
2.4 Study links
The relations table captures cross-study relationships that connect clinical trial records from different sources or define higher-order linkages such as reviews or citation networks. This table contains 940,582 relation instances and supports integration, traceability, and synthesis of trial-level evidence. Each row represents a directed edge between a head_id and a tail_id, with associated head_type and tail_type identifying the origin of the studies (e.g., PubMed, ClinicalTrials.gov, or regional registries). This structure allows the database to represent connections between registry entries and publications describing the same study, or group related records across different national databases. The semantic type of each link is encoded in the relation_type field, which includes labels such as linked_to (e.g., registry-publication pair), cite (e.g., PubMed-to-PubMed citation), include (e.g., a review including a trial), and exclude (e.g., a review explicitly excluding a study). These relationships enable downstream use cases such as study clustering, citation graph construction, and synthesis of trial networks for systematic reviews.
Figure 3: Manual data quality assessment across key database tables. The y-axis shows the proportion of valid (i.e., correctly extracted) records. Sample sizes (n) are shown below each bar.
2.5 Data quality assessment
To assess the quality of TrialPanorama, we employed seven annotators to manually verify approximately 1,500 records sampled across all major tables, stratified by data sources. Each sampled record was evaluated against the original source to determine its validity. As summarized in Figure 3, the database exhibits high overall accuracy, with most tables showing a validity rate above 90%. Notably, errors observed in some tables, particularly conditions and biomarkers, can be attributed to challenges in text-based pattern matching and the use of LLMs for information extraction. Despite these limitations, the results highlight the feasibility and promise of using LLMs to automatically structure clinical trial data from heterogeneous and unstructured sources.
3 TrialPanorama: Benchmark
TrialPanorama serves as a valuable resource for a wide range of clinical trial applications by offering a global view of trial protocols, outcomes, and systematic reviews. Building on this comprehensive foundation, we develop a suite of representative benchmark tasks spanning two key application areas: three tasks focused on systematic evidence synthesis and five tasks supporting data-driven clinical trial design. These benchmarks are illustrated in Figures 1b and 1c.
3.1 Systematic review tasks
The PRISMA statement [32] outlines standard procedures for conducting systematic reviews and meta-analyses in medicine. Among these, three critical steps are: (1) study search: identifying candidate studies from literature databases; (2) study screening: selecting eligible studies based on inclusion and exclusion criteria; and (3) evidence synthesis: aggregating findings from included studies to answer a targeted clinical question, such as the efficacy of a therapy for a specific condition. Based on TrialPanorama, we construct benchmark tasks corresponding to each of these key steps: study search, study screening, and evidence summarization. By leveraging the relations table, we are able to identify linkages between systematic review articles and the clinical trials they include or exclude, enabling the construction of high-quality, labeled benchmarks for these tasks. We hold the most recent studies and reviews out as the test set. The details of benchmark building, setups, and example samples are in Appendix B.
Study Search.
The study search task evaluates an AI model’s ability to retrieve relevant clinical trials from a large corpus, given the structured context of a systematic review. It simulates the first PRISMA step: identifying candidate studies based on a review’s topic. Each benchmark instance includes a query review defined by its background, objectives, and selection criteria, and models must retrieve trials that were included in the final review. Ground-truth labels are derived from the relations table in TrialPanorama. The benchmark includes 8,004 training reviews linked to 175,577 studies and 100 test reviews with 4,324 relevant studies. Retrieval performance is measured using Recall@K.
Study Screening.
The study screening task evaluates an AI model’s ability to determine whether candidate studies meet the inclusion criteria of a systematic review, simulating the second PRISMA step of study selection. Each benchmark instance includes a structured review query (background, objectives, and criteria) and 10 candidate studies described by both unstructured text (title and abstract) and structured metadata from TrialPanorama (e.g., phase, intervention, conditions). Ground-truth labels are derived from explicit inclusion/exclusion links in the relations table. Notably, the excluded candidates represent hard negatives, as they were manually excluded by the original review authors and recorded in the review’s study list. The benchmark includes 7,931 training reviews with over 175,000 included and 229,000 excluded studies, and 100 test reviews with 491 included and 488 excluded studies. Performance is evaluated using accuracy, precision, and recall.
Evidence Summarization.
The evidence summarization task evaluates an AI model’s ability to answer clinical research questions by synthesizing findings from relevant studies, emulating the last PRISMA step: evidence summarization. Each instance includes (1) the review’s structured setup (background, objectives, selection criteria), (2) a set of study abstracts containing relevant or irrelevant evidence, and (3) a multiple-choice question summarizing possible conclusions. Input data and gold labels are constructed using the relations and results tables in TrialPanorama, where study-review links define the context and structured trial findings inform the correct answer. Answer choices reflect typical clinical conclusions (e.g., “limited evidence” or “effective with adverse effects”). The benchmark contains 4,636 reviews and 35,512 studies in the training set, and 1,000 reviews with 8,974 studies in the test set. Model performance is assessed using accuracy and macro-F1 score.
3.2 Clinical trial design tasks
Designing a clinical trial involves multiple interdependent components that collectively determine its scientific validity and operational feasibility. As illustrated in Figure 1c, key design tasks include: (1) sample size estimation, which determines the number of participants needed to ensure sufficient statistical power to detect treatment effects; (2) eligibility criteria design, which defines the characteristics of the target population eligible for inclusion; (3) arm design, which specifies the allocation of participants into treatment and control groups with defined interventions; and (4) endpoint design, which identifies the clinical measures used to assess efficacy and safety. In addition, trial completion assessment evaluates whether a trial is likely to reach completion or be terminated early due to design flaws, recruitment challenges, or safety concerns, thus providing a holistic perspective on the robustness of the trial design. We hold the most recent studies out as the test set. The details of benchmark building, setups, and example samples are in Appendix C.
Arm, Eligibility Criteria, and Endpoint Design.
The trial design tasks assess a model’s ability to identify optimal design components for a given trial protocol. Each task presents a structured input, consisting of the trial’s title and brief summary, and asks the model to select the correct choice from a set of candidate design elements, where distractors are sampled from similar trials to ensure difficulty. In the arm design task, the goal is to select the correct configuration of intervention groups; in eligibility criteria design, the correct inclusion and exclusion conditions; and in endpoint design, the appropriate outcome measures. Candidate options are drawn from structured records in TrialPanorama, including data from the studies, dispositions, conditions, and endpoints tables. For all three tasks, performance is evaluated using classification accuracy. The benchmark includes 27,321 training studies and 1,000 test studies for each of the three tasks.
Sample Size Estimation.
The sample size estimation task evaluates a model’s ability to infer the required number of participants for a clinical trial, based on its structured protocol, including target conditions, interventions, and primary outcomes, under a fixed drop-out assumption. Each instance draws data from the studies, dispositions, and endpoints tables in TrialPanorama, paired with gold-standard target values extracted from protocols or publications. The benchmark includes 1,143 training examples and 1,000 test studies. Model performance is assessed using two metrics: accuracy (percentage of estimation within ±20% of ground truth) and mean absolute error (MAE). This task requires both clinical reasoning and quantitative estimation skills grounded in trial design logic.
Trial Completion Assessment.
This task evaluates a model’s ability to predict whether a clinical trial will reach completion or terminate early based on its protocol, and, if terminated, to identify the most likely reason (e.g., efficacy, safety, enrollment failure). Each instance includes structured trial metadata, such as study phase, arm count, masking, allocation, and eligibility criteria, drawn from the studies, dispositions, and results tables in TrialPanorama. The benchmark comprises 42,195 training studies and 1,000 test studies. Model performance is evaluated by balanced accuracy on (1) the binary completion prediction and (2) multiclass termination reason classification. This task reflects real-world feasibility assessments during trial planning.
4 Experiment
We selected a diverse set of state-of-the-art proprietary and open-source LLMs for evaluation. These include GPT-4o and its lightweight variant GPT-4o-mini, as well as the recent reasoning-focused model O3-mini, all accessed via the Azure OpenAI Service. In addition, we include the LLaMA-3.3-70B-Instruct and LLaMA-3.1-8B-Instruct models, representing large and moderate-scale open-source LLMs, accessed through the Azure AI Inference Service. This selection allows us to compare models across different scales, architectures, and reasoning capabilities.
Table 2: Benchmark experiment results for systematic review tasks.
Tasks Study search Study screening Evidence summarization
Models Recall@100 Recall@500 Accuracy Precision Recall Accuracy Macro-F1
Llama-8B 11.6 16.0 55.3 46.8 58.2 70.4 71.1
Llama-70B 25.1 33.3 60.7 59.1 70.5 78.8 76.7
GPT-4o-mini 17.3 22.7 60.2 60.4 60.5 74.1 73.1
GPT-4o 22.9 27.3 66.1 64.4 72.3 72.8 65.2
O3-mini 27.6 37.8 71.6 71.9 71.3 72.7 53.8
4.1 Systematic review tasks
Table 2 presents the performance of various LLMs across three core systematic review tasks: study search, study screening, and evidence summarization. The results highlight both general limitations and comparative strengths across models. First, performance on the study search task is uniformly low, with Recall@100 ranging from 11.6 (LLaMA-8B) to 27.6 (O3-mini). This reflects a key weakness of general-purpose LLMs: they struggle to generate precise and discriminative search queries that can retrieve relevant trials from a large corpus. Notably, O3-mini outperforms both proprietary (GPT-4o, GPT-4o-mini) and open-source (LLaMA-70B, LLaMA-8B) models by a significant margin, suggesting its reasoning-focused training may better align with the query formulation challenge.
Second, study screening performance is moderate, with accuracy scores ranging from 55.3 (LLaMA-8B) to 71.6 (O3-mini). The task is intrinsically difficult, as the negative examples in the test set are hard negatives. These studies are closely related to the review but were explicitly excluded by the original systematic review authors. Despite this challenge, O3-mini demonstrates the strongest overall performance in accuracy, precision, and recall, followed by GPT-4o. LLaMA-70B shows reasonable performance, while LLaMA-8B significantly underperforms.
Third, in evidence summarization, model performance varies more widely. LLaMA-70B achieves the highest accuracy (78.8) and Macro-F1 (76.7), outperforming GPT-4o and GPT-4o-mini, which show moderately strong but less consistent results. Surprisingly, O3-mini, which performs best on study search and screening, shows notably weak performance in summarization (Macro-F1 = 53.8), despite being positioned as a reasoning-oriented model. This result is somewhat counterintuitive, as reasoning-capable models are typically expected to excel at inference tasks that involve aggregating evidence from multiple sources.
Table 3: Benchmark experiment results for clinical trial design tasks. “Acc”: short for accuracy; “BAcc-out”: short for balanced accuracy of trial outcome assessment results; “BAcc-rt”: short for balanced accuracy of trial termination rationalization results; “MAE $\downarrow$”: mean absolute error that is the smaller the better.
Tasks Arm design Eligiblity criteria design Endpoint design Sample size estimation Trial completion
Models Acc Macro-F1 Acc Macro-F1 Acc Macro-F1 Acc MAE $\downarrow$ BAcc-out BAcc-rt
Llama-8B 79.9 79.9 55.0 54.6 52.1 52.9 3.3 1760.6 54.0 15.7
Llama-70B 84.6 84.7 72.6 73.1 61.8 62.5 18.7 633.3 52.6 17.9
GPT-4o-mini 78.4 78.7 55.8 55.8 62.2 62.2 22.8 411.4 51.3 16.8
GPT-4o 85.0 85.0 76.3 76.5 69.1 69.4 22.9 512.6 59.3 19.5
O3-mini 85.9 86.1 80.4 80.6 52.1 60.7 25.2 452.1 51.0 17.0
4.2 Clinical trial design tasks
Table 3 presents the evaluation of LLMs across five clinical trial design tasks. First, arm design appears to be the easiest among the structured design tasks, with all models achieving relatively high accuracy and macro-F1. O3-mini achieves the highest scores (85.9 accuracy and 86.1 macro-F1), followed closely by GPT-4o, suggesting that LLMs are generally capable of recovering structured intervention information when explicitly described in the trial protocol.
In contrast, endpoint design proves to be more challenging. All models perform worse on this task, with accuracy ranging from 52.1 to 69.1 and lower macro-F1 across the board. The relatively poor performance, even among advanced models like GPT-4o and O3-min. Second, sample size estimation remains a particularly difficult task for all models. Despite the reasoning-oriented nature of the problem, no model achieves an accuracy above 26%, and mean absolute error (MAE) remains high for all (e.g., 411.4 for GPT-4o-mini and 452.1 for O3-mini). This indicates that LLMs, even those designed for reasoning, lack the statistical grounding or calibration mechanisms required to estimate participant counts from trial parameters such as design type, treatment arms, and expected effect size. Third, trial completion assessment is also highly challenging. Performance on binary outcome prediction (whether a trial is completed or terminated) and termination rationale identification appears near random, with balanced accuracy scores clustered around 50–55% for trial status (BAcc-out) and 15–20% for termination reasoning (BAcc-rt). This reflects the difficulty of inferring future feasibility outcomes from trial protocols.
In a nutshell, these results indicate that current general-purpose LLMs fall short in supporting key trial design decisions. While they show some promise in recognizing explicit design elements, they are far less reliable for tasks that require statistical reasoning, feasibility forecasting, or context-aware endpoint selection. These gaps highlight the need for further research and specialized model development for clinical trial tasks.
5 Related work
Several recent efforts have focused on constructing structured datasets and benchmarks for clinical trials. TOP [19] introduced a benchmark for clinical trial outcome prediction with approximately 10K trials, while TrialBench [30] expanded to include 400K trials from ClinicalTrials.gov for predicting trial duration, outcomes, and other attributes. CTKG [11] proposed a knowledge graph that captures design entities (e.g., drugs, conditions) from 8K trials, and FinegrainedResults [9] structured 56K trials into arm-level result graphs. CTO [14] similarly focused on representing outcome data for 125K interventional trials. CARE [10] extracted structured trial results from 700 literature publications, while PubMedKG [31] linked trial identifiers and citations across publications, patents, and trial registries. Other lines of work have developed benchmarks for systematic literature review (SLR) tasks. TrialReviewBench [28] built a benchmark with 100 systematic reviews for study search, screening, and evidence summarization, and LEADSInstruct [27] scaled these efforts to 21K reviews. However, most of these datasets are limited in scope, focusing only on trials’ design or results, lacking ontology alignment, or without accompanying benchmark tasks. Our work, TrialPanorama, addresses these limitations by unifying trial records from 16 global sources and linking them to biomedical ontologies. It supports both structured design and result representations, and enables a broad suite of benchmark tasks for both trial design and systematic review automation.
6 Discussion and conclusion
Despite the broad coverage and structured integration of clinical trial data in TrialPanorama, several limitations should be acknowledged. First, the dataset may reflect biases introduced by the data collection process itself. For instance, trials registered on ClinicalTrials.gov, primarily based in the United States, are more consistently structured and accessible than those from other countries, potentially leading to overrepresentation of U.S.-centric studies or specific trial designs. Similarly, trial information extracted from scientific publications relies on unstructured text, which varies in quality and completeness and may introduce processing bias. Second, there is inherent selection bias in the publication landscape, where studies with positive outcomes are more likely to be published, skewing the dataset’s representation of trial attributes and results. Third, while LLMs were employed to assist in data extraction and have demonstrated promising accuracy based on sampled validation, systematic evaluation and error mitigation remain areas for future improvement to ensure high data fidelity across the entire dataset. Lastly, the benchmark data constructed in this work represents a simplified and task-focused abstraction of real-world clinical trial workflows. It does not capture the full complexity and variability of trial design and execution. In addition, the selected LLMs and baseline methods evaluated here reflect only a subset of possible approaches; further exploration of model variants, the role of memorization, and prospective evaluation settings will be important for understanding generalizability and real-world utility.
To summarize, we presented TrialPanorama, a large-scale structured database of clinical trials, along with a benchmark of eight tasks for systematic review and trial design. Evaluation of five leading LLMs reveals limited performance, highlighting the need for domain-adapted AI systems in clinical research. TrialPanorama provides a foundation to support future development of intelligent tools for clinical trial planning, evidence synthesis, and automation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: [N/A]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: [N/A]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not involve theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: [N/A]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: [N/A]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: [N/A]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: [N/A]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: [N/A]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: [N/A]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: [N/A]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: [N/A]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
[…]
1Contents of Appendix
Appendix A Database building
A.1 Raw data collection
To construct our benchmark dataset, we crawled clinical trial and publication records from a diverse set of publicly accessible registries and databases. The data URLs and the date of the most recent crawl are listed in Table 4.
Table 4: List of data sources and the corresponding last crawl dates.
Source Data URL Last Crawled
ANZCTR https://www.anzctr.org.au/ April 1, 2025
Brazilian Registry https://ensaiosclinicos.gov.br April 1, 2025
ChiCTR https://www.chictr.org.cn/searchprojEN.html April 1, 2025
Dutch Registry https://www.onderzoekmetmensen.nl/en/ April 1, 2025
EU Clinical Trials https://www.clinicaltrialsregister.eu April 1, 2025
DRKS https://drks.de/search/en/results April 1, 2025
Iranian Registry https://irct.behdasht.gov.ir/ April 1, 2025
ISRCTN https://www.isrctn.com April 1, 2025
Japan Registry https://center6.umin.ac.jp/cgi-open-bin/ctr_e/ctr_view.cgi April 1, 2025
South Korea CRIS https://cris.nih.go.kr/cris/index/index.do April 8, 2025
Pan African Registry https://pactr.samrc.ac.za/Search_v2.aspx April 8, 2025
Sri Lanka CTR https://slctr.lk/trials April 8, 2025
Thai Registry https://www.thaiclinicaltrials.org April 8, 2025
REec https://reec.aemps.es/reec/public/web.html April 9, 2025
CTRI https://ctri.nic.in/Clinicaltrials/advancesearchmain.php April 9, 2025
PubMed https://ftp.ncbi.nlm.nih.gov/ April 9, 2025
After collection, the raw data was subjected to systematic cleaning procedures to ensure consistency and reliability across sources:
  • Intra-source de-duplication: Duplicate entries within the same registry were detected and removed using unique identifiers (such as trial registration numbers), and corroborated using additional metadata like titles.
  • Inter-source de-duplication: Clinical trials registered in multiple registries were identified via fuzzy matching on titles. Duplicate records across sources were merged into a single one.
  • Field normalization: Metadata fields such as trial status, registration date, and condition labels were standardized across sources to ensure consistent querying.
  • Missing value handling: Entries with missing or incomplete critical fields were flagged.
These operations minimized noise, reduced redundancy, and created a unified and clean dataset suitable for benchmarking tasks.
A.2 Studies table
There are three main sources of clinical trial records in our dataset: ClinicalTrials.gov, global registries, and PubMed. ClinicalTrials.gov records are highly structured, and we extracted key attributes such as trial type, sponsor, start year, and phase from the AACT database  [^1] https://aact.ctti-clinicaltrials.org/.
In contrast, trial data from global registries are more heterogeneous and often semi-structured. Taking the processing of ANZCTR records as an example, we manually curated and standardized each trial entry to align with our unified schema. First, we used the "Registration number" column as the unique identifier for each study and assigned it as the ‘study_id‘. Trial results were extracted from the "Results – plain English summary" field and, when available, supplemented with citation links from "Citation/DOI/link/details." We also processed structured eligibility information: for instance, the gender field was standardized by mapping values like "Both males and females" to a normalized label "MALE/FEMALE," ensuring consistency across sources. Age eligibility was retained from the "Minimum age" and "Maximum age" fields, and healthy volunteer status was preserved from the corresponding binary field. Trial phases were parsed by checking substrings within the "Phase" column and converting them into discrete labels such as ‘PHASE1‘, ‘PHASE2‘, etc. Additionally, we extracted outcome information from the "Primary outcome" and "Secondary outcome" fields and abstracted trial objectives from the “Brief summary” column. These extracted attributes were compiled into a standardized table that serves as the ANZCTR subset of the merged dataset.
Moreover, clinical trials from publications like PubMed are essentially unstructured, typically presented as free-text abstracts without standardized fields. To extract structured information, we developed a rule-based pipeline that leverages regular expressions to identify key attributes embedded in the abstract. For each paper, we first extract the PubMed ID as the unique ‘study_id‘ and retain the abstract and title as primary textual content. We then parse trial phases by searching for standard mentions such as “Phase I” or “Phase 3b” across both title and abstract, and map these to a normalized phase label (e.g., ‘PHASE1‘, ‘PHASE2‘, etc.). To approximate sample size, we detect phrases like “XX participants” or “XX patients” and extract the numerical value as the actual accrual count. Additionally, we identify trial results by aggregating sentences prefixed with sections such as “Results:” or “Conclusions:” when available in the abstract. This approach allows us to recover basic outcome information despite the absence of structured metadata. While this method does not capture the full complexity of trial protocols, it provides a scalable way to convert a large volume of unstructured publications into structured records suitable for downstream analysis.
A.3 Conditions table
To standardize disease conditions associated with each trial, we extracted Medical Subject Headings (MeSH) terms using the NCBO BioPortal Annotator API [^2] https://bioportal.bioontology.org/annotator. This process was applied only to non–ClinicalTrials.gov sources (e.g., PubMed, Embase, and international registries), as ClinicalTrials.gov conditions are already standardized. We used the official title and brief summary of each trial as input to retrieve disease-related MeSH terms (semantic type T047) and their hierarchical ancestors. The resulting MeSH types were labeled as mesh-list (primary condition), mesh-ancestor (intermediate ancestors), and mesh-ancestor-main (top-level ancestor).
A.4 Drugs Table
To construct a comprehensive drugs table, we extract intervention mentions from three primary sources: (1) ClinicalTrials.gov and (2) PubMed. For ClinicalTrials.gov, we include entries where intervention_type is labeled as DRUG. For PubMed, we extract the intervention component (“I”) from PICO triplets and apply a lexicon-based filter to retain only drug-related mentions, excluding non-drug interventions such as procedures, devices, or behavioral treatments.
All extracted raw mentions are then processed through a multi-stage drug normalization and mapping pipeline (Figure 1). The pipeline begins with name normalization: (1) removing common treatment-related suffixes such as “therapy” or “tablet”, (2) cleaning noisy strings by removing typographic artifacts, and (3) handling special cases including abbreviations and regional variants.
Next, drug names undergo RxNorm resolution via a dual approach combining local and remote resources. (1) Local mapping attempts to match names against a pre-indexed RxNorm database with curated mappings. (2) If local matching fails, we query the official RxNorm API, parse the returned XML, and extract the corresponding RXCUI. (3) If both methods fail, we request spelling suggestions from the API and retry mapping using the suggested variants.
For drug names successfully mapped to RxNorm, we proceed to DrugBank resolution. (1) We first attempt to map the resolved RXCUI to a DrugBank ID via a curated RxNorm-to-DrugBank mapping table. (2) If that fails, or if RxNorm mapping was unsuccessful, we apply a direct DrugBank matching routine using exact string matching, synonym expansion, brand name resolution, and fuzzy matching based on character edit distance.
This process results in three possible outcomes: (1) drug names mapped to both RxNorm and DrugBank, (2) drug names mapped to DrugBank only, and (3) drug names that remain unmapped after all resolution steps.
Table 5 summarizes the drug mapping statistics for each data source, including total unique drug mentions, RxNorm mapping counts and rates, and DrugBank resolution results.
Table 5: Drug mapping statistics across data sources.
Source Total Drugs RxNorm Mapped RxNorm Rate (%) DrugBank Mapped DrugBank Rate (%)
PubMed 953,361 579,035 60.74 241,879 25.37
ClinicalTrials.gov 366,214 290,371 79.29 273,509 74.69
Supplementary Figure 1: Pipeline for drug name normalization and standardization. Drug names are cleaned, mapped to RxNorm, and subsequently resolved to DrugBank identifiers using a combination of indirect and direct strategies.
To further enrich the standardized drug table, we annotate each drug entry with regulatory approval status. Specifically, we compile lists of drugs approved by three major regulatory agencies: the U.S. Food and Drug Administration (FDA), the European Medicines Agency (EMA), and Japan’s Pharmaceuticals and Medical Devices Agency (PMDA). These lists are sourced from DrugCentral 2023 [33] [^3] https://drugcentral.org/, which provides a consolidated view of approval data across jurisdictions.
Approval annotations are applied by checking each drug’s normalized name, RxNorm-mapped name, and DrugBank-mapped name against the cleaned lists of approved drug names. A drug is marked as approved by a given agency if any of these name variants match an entry in the corresponding list. This process enables us to systematically track the clinical regulatory status of drugs within our dataset, aiding downstream tasks that require distinguishing between approved and experimental interventions.
A.5 Trial results table
To construct the Trial Result Table, a core component for the evidence summarization benchmark task, we systematically extracted key findings from two primary types of sources within our TrialPanorama database: 1) published clinical trial literature, encompassing abstracts from PubMed, and 2) structured trial data directly from ClinicalTrials.gov registry entries. For all sources, the extraction process focused on identifying the core PICO (Population, Intervention, Comparator, Outcome) elements.
For published trial literature, a large language model (LLM), gpt-4o-mini, automatically extracted PICO components from PubMed abstracts. For data from ClinicalTrials.gov, a custom rule-based parser processed the structured JSON trial records, focusing on those with reported results (i.e., resultsSection present and outcome measures having a reportingStatus of ’POSTED’). The extracted trial results include:
  • P (Patient or problem): Study population, conditions, and characteristics.
  • N (Number of subjects): Total participants in the population group.
  • Outcomes (per arm): For each reported outcome, the associated intervention arm (I), its status as a control (C, boolean flag), and detailed outcome results (O), including quantitative data and conclusions.
This process yielded a consistent, structured output for each abstract and document from ClinicalTrials.gov. Information extracted from both published literature and ClinicalTrials.gov was harmonized and consolidated into a unified Trial Result Table. Each row in this table represents a specific outcome from a particular trial arm. The prompt for LLM extraction is shown below.
Given the following clinical trial paper abstract, your task is to extract the key trial information using the PICO framework. The PICO framework consists of the following components:
Patient or problem (P): Who is the patient or what is the population or condition being addressed? Report detailed information including the population size, condition, and any relevant characteristics.
Intervention or control (I/C): What is the main intervention, treatment, or control being considered? Report detailed information including the intervention, including dosage and duration if applicable.
Outcome(s) (O): Extract detailed results and conclusions related to the outcomes, including numbers, p-values, and outcome-related and statistical conclusions if applicable.
Please return a list of reported outcome measures. Each reported outcome item should include the intervention (I) or Control (C) used and the detailed corresponding outcomes (O) of the intervention / control. Flag the control group if applicable. Only extract the information that is explicitly mentioned in the abstract. If any of the components are not mentioned in the abstract, please indicate that they are not available.
Title: {title}
Abstract: {abstract}
A.6 Trial outcomes table
Supplementary Figure 2: (a) Trial outcome extraction workflow. Completed trials are labeled using outcome measures and matched PubMed abstracts. Terminated trials are categorized by mapping stop reasons to predefined outcome types via keyword similarity. (b) Distribution of extracted outcome types on trials from ClinicalTrials.gov.
To extract trial outcomes, we developed two distinct workflows for trials registered on ClinicalTrials.gov and those from other international sources. For ClinicalTrials.gov trials, we first categorized them based on overall status into: (1) completed, (2) terminated/withdrawn/suspended, and (3) other (e.g., recruiting, not active, withheld). We focused on completed and terminated trials to derive fine-grained outcome types using multiple sources, including publications and reported results, as shown in Supplementary Figure 2.
For completed trials, if explicitly labeled as approved for marketing, this outcome type was directly recorded. For the remaining trials, we identified matched publications by selecting the top two PubMed abstracts most relevant to the trial’s official title using text similarity. We also extracted primary outcome measures from ClinicalTrials.gov. To determine whether a trial had a positive or negative outcome, we used GPT-4o to infer if the primary endpoint was met, based on the trial title, matched abstracts, and outcome measures. For trials without relevant abstracts or unclear outcomes, we used reported p-values for primary outcomes, applying two thresholds (0.01 and 0.05) to assign outcome types. The prompt for outcome type extraction as follows:
TASK:
You are a clinical trial expert. You are given a clinical trial with NCT ID [nct_id] and its official title: [official_title].
The primary outcomes of the trial are: [primary_outcomes].
Based on the related articles of the trial given below, please determine if the trial is likely to have a positive or negative outcome.
A positive outcome means that the trial has achieved any of its primary endpoints or outcomes. Negative outcome means that the trial has not achieved any of its primary endpoints or outcomes.
If the outcome is unclear or unable to be inferred from the articles below, please choose "unknown".
Articles:
1. Title: [top_1_similar_article_title]
Abstract: [top_1_similar_article_abstract]
1. Title: [top_2_similar_article_title]
Abstract: [top_2_similar_article_abstract]
OUTPUT:
Please provide a short answer by choosing one of the following options:
positive outcome
negative outcome
unknown
Do not provide any explanation or additional information.
To derive fine-grained outcome types for terminated trials, we first embedded the stop reasons and visualized their clusters to identify representative keywords, which were grouped into predefined outcome types (Supplementary Figure 2). We further enriched these keywords using an LLM. For each trial, we computed similarity scores between the reported reason and the enriched keyword sets using a cross-encoder (ms-marco-MiniLM-L-12-v2), and assigned the outcome type with the highest average similarity. For trials from non-U.S. registries, we filtered those with reported results, which were in unstructured text format. We provided the trial title, primary outcome measures, and result text to GPT-4o to infer the trial outcome.
A.7 Relations table
Table 6: Statistics of the relation types in the relations table.
Relation Type Count Percentage (%)
PubMed - cite - PubMed 349,680 37.177
PubMed - exclude - PubMed 237,076 25.205
PubMed - include - PubMed 180,827 19.225
PubMed - linked_to - ClinicalTrials.gov 167,121 17.768
ClinicalTrials.gov - linked_to - JAPAN 2,787 0.296
PubMed - include - ClinicalTrials.gov 954 0.101
ClinicalTrials.gov - linked_to - KOREA 632 0.067
PubMed - cite - ClinicalTrials.gov 425 0.045
PubMed - exclude - ClinicalTrials.gov 406 0.043
ClinicalTrials.gov - linked_to - GERMAN 331 0.035
ClinicalTrials.gov - linked_to - ANZCTR 201 0.021
ClinicalTrials.gov - linked_to - PAN_AFRICAN 56 0.006
ClinicalTrials.gov - linked_to - SRI_LANKA 37 0.004
ClinicalTrials.gov - linked_to - CHICTR 31 0.003
ClinicalTrials.gov - linked_to - DUTCH 18 0.002
We implemented a multi-stage data processing pipeline centered on Cochrane Library reviews and their linked PubMed metadata to construct structured relationships between systematic reviews, clinical trials, and associated publications. Each review was first matched to its corresponding PubMed ID, and the full-text XML was parsed to extract structured reference sections. Using a custom reference extraction function, we identified citations grouped under section titles such as references to studies included in this review, excluded from this review, and awaiting assessment. For each cited study, we extracted the PubMed ID and, when available, trial registry identifiers such as NCT numbers using regular expressions applied to the citation text. We formalized these relationships as triples in the format (head_id, relation_type, tail_id), where head_id is the PubMed ID of the review, tail_id is the PubMed ID of the referenced study, and relation_type is one of include, exclude, or cite, denoting the review’s assessment of the study.
We further enriched these relationships by linking PubMed studies to ClinicalTrials.gov trial records. This linkage was established through two mechanisms: first, by inspecting the AccessionNumber field in the PubMed XML to identify embedded NCT numbers; and second, by applying regular expressions to the abstract text to extract any mentioned NCT identifiers. In addition, we extended this mapping to include cross-registry linkages. For example, NCT numbers occasionally appear in records from other national trial registries; in such cases, we created edges of type link_to to represent the equivalence between records referring to the same underlying trial across different sources. The statistics of all the relation types are in Supplementary Table 6.
A.8 Dispositions table
Disposition reflects an essential aspect of a clinical trial: the design of experimental arms, including details such as the target treatment, administration schedule, duration, and dosage. To extract this information from ClinicalTrials.gov, we processed two complementary data sources: the trial design groups and their associated interventions. We began by loading and standardizing the relevant tab-separated files, namely design_groups, interventions, and the linking table design_group_interventions. Key columns were normalized (e.g., renaming id to design_group_id or intervention_id), and the data were merged using shared nct_id and identifier fields. This enabled us to associate each design group with its corresponding intervention metadata. The resulting dataset provides, for each trial arm, a structured mapping that includes the intervention type (e.g., drug, device), group type (e.g., experimental, active comparator), treatment name, and descriptive text detailing the treatment protocol.
For publications, we extracted trial results in PICO format, where each unique pairing of the Population (P) and Intervention (I) elements is treated as a distinct disposition. From these pairs, we constructed corresponding trial arms that align with the disposition structure used in ClinicalTrials.gov. Each derived arm includes fields such as treatment name, description, and group type, allowing for consistent representation and integration across both registry-based and publication-based sources.
A.9 Biomarkers table
Supplementary Figure 3: The distribution of the most frequent biomarkers under each of the five therapeutic biomarker classes.
Biomarkers, particularly therapeutic biomarkers (ThMARs), are critical to the design and execution of clinical trials. ThMARs have become essential in both clinical development and practice across therapeutic areas. Five major types of ThMARs have been identified, each playing a distinct role in drug discovery and development: (1) Pharmacodynamic Biomarkers, which ensure the pharmacological effect of a therapy; (2) Safety Biomarkers, used to assess the likelihood or severity of therapy-induced toxicity; (3) Monitoring Biomarkers, which guide ongoing clinical management by tracking patient status over time; (4) Predictive Biomarkers, which help tailor therapies to individuals most likely to benefit; and (5) Surrogate Endpoints, which serve as proxies to accelerate therapy approval.
We leveraged TheMarker [34] as a curated reference source for therapeutic biomarkers (ThMARs) relevant to clinical trials. Our processing pipeline to identify and normalize biomarker mentions consisted of two stages. In the first stage, we employed a GPT-4o-mini model to extract candidate text spans from clinical trial descriptions, such as titles, abstracts, and outcome sections, that may contain therapeutic biomarkers. The model was prompted to directly extract potentially relevant mentions verbatim from the source text. For example, given a trial abstract, the model might identify spans such as “HER2-positive breast cancer” or “PD-L1 expression level” as candidate biomarkers.
In the second stage, we implemented a deterministic pattern matching module to normalize these candidate spans by linking them to standardized biomarker entries from TheMarker. This was achieved through a two-step matching strategy. First, we attempted an exact string match between the extracted phrase and the entries in a pre-indexed lookup table. If an exact match was found, the matched biomarker ID and corresponding metadata were recorded, and the phrase was classified with the match type exact_multi_word. If no exact match was found, we then attempted an order-invariant match by sorting the tokens in the candidate phrase and comparing them against a token-sorted index of biomarker terms. This allowed us to recover phrases with slight reordering of words, such as matching “receptor estrogen alpha” to “estrogen receptor alpha,” and was labeled as exact_multi_word_order_invariant.
Each successful match was associated with a structured biomarker representation that includes: (1) biomarker_name, the canonical name of the matched biomarker; (2) biomarker_class, describing the functional or molecular category of the biomarker (e.g., protein, receptor); (3) biomarker_type, which corresponds to one of the five therapeutic biomarker types defined by TheMarker: Pharmacodynamic, Safety, Monitoring, Predictive, or Surrogate Endpoint; and (4) biomarker_genes, listing any specific genes associated with the biomarker when applicable. This structured mapping enables precise linkage between clinical trials and biomarker knowledge, facilitating downstream analyses on biomarker-driven trial design and therapeutic targeting.
Supplementary Figure 3 shows the distribution of the most frequent biomarkers under each of the five therapeutic biomarker classes. In the Predictive (PRD) category, C reactive protein and Tumor necrosis factor appear most frequently, reflecting their widespread use in identifying responsive patient subpopulations. The Surrogate (SUR) category is heavily dominated by Blood pressure, with over 60,000 mentions, followed by metrics such as Bone mineral density and Serum creatinine, commonly used as clinical endpoints in chronic disease trials. In the Pharmacodynamic (PDY) class, Low density lipoprotein and Angiotensin converting enzyme are among the most prominent, reflecting their roles in evaluating pharmacological activity, particularly in cardiovascular studies. For the Safety (SAF) biomarkers, Granulocyte macrophage colony-stimulating factor and Hyaluronic acid are the most frequently referenced, indicating their importance in monitoring adverse or toxic effects. Lastly, in the Monitoring (MOI) category, Epidermal growth factor, HBV DNA, and CEA (Carcinoembryonic antigen) are among the top markers used for longitudinal assessment of disease progression or therapeutic response. Overall, this distribution highlights the heterogeneity in biomarker utility across different therapeutic contexts, with some biomarkers, such as C reactive protein and Blood pressure, being widely applicable across multiple trial types.
A.10 Endpoints table
Core Outcome Measurement Sets [35, 36] provide a standardized taxonomy for classifying clinical trial outcomes, comprising five high-level domains and their associated subdomains. We adopted this COMET taxonomy to categorize outcome measurements reported in clinical trials. To automate the classification process, we employed an in-context learning approach using a GPT-4o-mini model. For each trial, the model was prompted with structured instructions to first extract the primary outcome measurement from the trial abstract or description, and then assign it to the appropriate COMET domain and subdomain. The full prompt used in this process is shown below.
You are a clinical research expert. Given a clinical trial abstract or description, extract the PRIMARY OUTCOME MEASUREMENT (usually only one or at most three), and classify them using the COMET taxonomy.
Follow these steps:
1. Extract the PRIMARY OUTCOME MEASUREMENT from the input text.
2. Classify the PRIMARY OUTCOME MEASUREMENT using the COMET taxonomy below.
COMET Taxonomy:
1. Mortality/survival
- No subdomains
2. Physiological/clinical
- Blood and lymphatic system outcomes
- Cardiovascular outcomes
- Endocrine outcomes
- Gastrointestinal outcomes
- General outcomes
- Immune system outcomes
- Infections and infestations outcomes
- Metabolic and nutritional outcomes
- Musculoskeletal and connective tissue outcomes
- Neoplasms and related outcomes
- Nervous system outcomes
- Pregnancy and childbirth outcomes
- Renal and urinary outcomes
- Reproductive system and breast outcomes
- Respiratory outcomes
- Skin and subcutaneous tissue outcomes
- Surgical and medical procedures outcomes
- Vascular outcomes
3. Life impact
- Delivery of care
- Global quality of life
- Perceived health status
- Personal circumstances
- Physical functioning
- Role functioning
- Social functioning
- Emotional functioning/wellbeing
- Cognitive functioning
- Sleep
- Fatigue
4. Resource use
- Hospital
- Need for further intervention
- Societal/carer burden
- Other resource use
5. Adverse events/effects
- Any subdomain
Return your output as a list of JSON objects, one per outcome, in the following format:
"outcome": "<brief name of the primary outcome measurement>",
"domain": "<COMET domain>",
"subdomain": "<COMET subdomain or null>"
Input text:
{outcome_text}
The prompt included the full taxonomy definitions as context, covering domains such as Mortality/Survival, Physiological/Clinical, Life Impact, Resource Use, and Adverse Events, along with their respective subdomains. The model output was formatted as a list of JSON objects, each containing the outcome name, its mapped domain, and subdomain. This approach allowed for flexible, semantically aware interpretation of unstructured outcome descriptions, reducing the need for hand-crafted rules while achieving fine-grained alignment with the COMET classification. The result is a structured, interpretable dataset of standardized clinical trial endpoints that facilitates cross-trial comparisons, systematic reviews, and meta-analyses.
A.11 Adverse events table
The Adverse Event Table is systematically populated by processing the adverseEventsModule found within the resultsSection of structured JSON records from ClinicalTrials.gov. The primary aim of this table is to provide a comprehensive overview of the safety profiles reported in these clinical trials.
Adverse event information was extracted from ClinicalTrials.gov data using a custom rule-based parser. We process the JSON records to identify and structure reported adverse events. The key information extracted includes:
  • Trial Arm Identification (I/C): Distinct trial arms (or "event groups") are identified from the eventGroups list. The title and description of each group are combined to form a descriptive string for the intervention or comparator arm.
  • Adverse Event Details: For each arm, both "serious" (seriousEvents) and "other" (otherEvents) adverse event categories are processed. For each reported event, the parser extracts:
    • The specific term used to describe the event (e.g., "Nausea").
    • The organSystem affected.
    • The frequency of the event within that specific arm, detailed as numAffected (number of participants experiencing the event) and numAtRisk (total participants in that arm). An event is only recorded for an arm if numAffected is greater than zero.
  • Structured Data per Event Instance: For each adverse event occurring in a specific trial arm, the following structured information is captured:
    • The adverse event Term.
    • A comprehensive Description string that includes the organSystem, any associated notes, and the specific frequency of the event in that arm (formatted as "Frequency: X affected / Y at risk").
    • An is_serious boolean flag to indicate if the event was classified as serious by the trial reporters.
The extracted and structured adverse event data is then assembled into a flattened tabular format. Each row in the final Adverse Event Table represents a single adverse event term as it was reported for a specific trial arm.
To enhance standardization and facilitate analysis, the reported adverse event Terms are mapped to the Medical Dictionary for Regulatory Activities (MedDRA). This mapping utilizes the MedDRA hierarchy files (LLT, PT, HLT, HLGT). The corresponding MedDRA code is then added to the table.
This systematically constructed Adverse Event Table, linked by NCTID and trial arm descriptions, serves as a valuable resource for safety analyses and complements the efficacy-focused Trial Result Table within the TrialPanorama database.
Appendix B Benchmark building: Systematic review tasks
B.1 Study search
Table 7: Example instance of Study Search
Source PubMed ID: 38078494 Question: Generate search query to find PubMed studies based on the systematic review setup as below. Background: Epilepsy is a common neurological disorder, affecting 0.5% to 1% of the population. In nearly 30% of cases, epilepsy is resistant to currently available drugs. Pharmacological treatment remains the first choice to control epilepsy. Lamotrigine is a second-generation antiseizure medication. When used as an add-on (in combination with other antiseizure medications), lamotrigine can reduce seizures, but with some adverse effects.", Objectives: "To evaluate the benefits and harms of add-on lamotrigine, compared with add-on placebo or no add-on treatment in people with drug-resistant focal epilepsy.", Selection criteria: We included randomised controlled trials (RCTs) that investigated add-on lamotrigine versus add-on placebo or no add-on treatment in people of any age with drug-resistant focal epilepsy. We used data from the first period of eligible cross-over trials. Ground-truth studies (PubMed ID) "20696552", "2612495", "8937535", "10563619", "2498073", "2127016", "8232944", "8112232", "17938371", "18077797", "8505632", "8453943"
Definition & Motivation.
This task simulates the study search step in conducting systematic reviews, where the model must comprehend a given research question and synthesize a corresponding search query to retrieve relevant studies for inclusion. The key motivation behind this task is to streamline a traditionally labor-intensive process that typically requires the expertise of trained librarians, particularly due to the complexity and variability of search query formats across literature databases. Automating this step can significantly accelerate systematic review workflows. An example instance is presented in Supplementary Table 7.
Input & Output.
The input for this task includes the setup of a systematic review, consisting of the research background, objective, and study selection criteria. The model is expected to generate a structured search query, which is then submitted to the PubMed Search API to retrieve the top 500 results and their associated PubMed IDs. A successful output should go beyond simple keyword extraction, incorporating semantic understanding and augmentation of the input to produce a comprehensive and effective query.
Dataset Construction.
We constructed the dataset using a relation table that links systematic reviews to primary studies. Specifically, we focused on instances with the relation type include, indicating that the study was selected for inclusion in the review. For each review, we parsed its full text to extract relevant sections and used the linked PubMed abstracts as ground-truth targets. This process resulted in a dataset of 8,104 examples derived from systematic review articles, covering more than 180,000 included PubMed studies. We reserved the most recent 100 reviews as a held-out test set.
Evaluation Metrics.
We evaluate this task as a retrieval problem by measuring how many ground-truth studies appear in the top search results produced by the model-generated query. Specifically, we report Recall@100 and Recall@500, reflecting the number of included studies retrieved among the top 100 and 500 PubMed results, respectively.
Table 8: Example instance of Study Screening
Source: PubMed ID: 38994714 Background: Cystic echinococcosis is a parasitic infection mainly impacting people living in low- and middle-income countries. Infection may lead to cyst development within organs, pain, non-specific symptoms or complications including abscesses and cyst rupture. Treatment can be difficult and varies by country. Treatments include oral medication, percutaneous techniques and surgery. One Cochrane review previously assessed the benefits and harms of percutaneous treatment compared with other treatments. However, evidence for oral medication, percutaneous techniques and surgery in specific cyst stages has not been systematically investigated and the optimal choice remains uncertain. Objective: To assess the benefits and harms of medication, percutaneous and surgical interventions for treating uncomplicated hepatic cystic echinococcosis. Selection Criteria: We included randomized controlled trials (RCTs) in people with a diagnosis of uncomplicated hepatic cystic echinococcosis of World Health Organization (WHO) cyst stage CE1, CE2, CE3a or CE3b comparing either oral medication (albendazole) to albendazole plus percutaneous interventions, or to surgery plus albendazole. Studies comparing praziquantel plus albendazole to albendazole alone prior to or following an invasive intervention (surgery or percutaneous treatment) were eligible for inclusion. Candidate Studies: PMID: 32382859 — Comparison of PAIR and catheterization techniques for CE1 and CE3a liver hydatid cysts. PMID: 35708380 — [Title and abstract truncated for brevity.] PMID: 16927868 — [Title and abstract truncated for brevity.] PMID: 24464258 — [Title and abstract truncated for brevity.] — (additional candidate studies omitted for space) Screening Labels: 32382859 — Included 35708380 — Included 16927868 — Excluded
B.2 Study screening
Definition & Motivation.
This task simulates the study selection step in a systematic review pipeline, where the model must assess the eligibility of candidate studies for inclusion based on the review’s predefined criteria. Given a systematic review’s background, objective, and eligibility criteria, along with a set of candidate studies retrieved from PubMed (each with a title and structured abstract), the model determines whether each study should be included or excluded. This step is often labor-intensive and requires detailed reasoning across clinical nuances and protocol designs. An example instance is shown in Supplementary Table 8.
Input & Output.
The input to this task consists of the systematic review protocol (comprising the review’s background, objective, and selection criteria) and a list of candidate studies, each represented by its title and structured abstract. The output is a binary label (included or excluded) for each candidate study, indicating whether it meets the review’s eligibility criteria. Accurate performance requires the model to match eligibility conditions against sometimes subtle or incomplete cues in the candidate study abstracts.
Dataset Construction.
We construct the dataset using reviews that have explicit inclusion and exclusion annotations for candidate studies, available from the relations table in our clinical trial knowledge base. For each review, we extract its background, objective, and selection criteria from the abstract, and retrieve associated candidate studies along with their labels (included or excluded) based on the relation types include and exclude. To ensure a balanced and informative evaluation, we downsample the candidate pool for each review to include a random mix of included and excluded studies, preserving diversity while controlling for class imbalance. Each instance contains up to 10 candidate studies, each with a title and structured abstract, paired with a single review protocol.
Evaluation Metrics.
We evaluate model performance using three metrics: accuracy, precision, and recall. Accuracy reflects the overall proportion of correctly classified studies. Precision measures the proportion of correctly identified included studies among all studies predicted as included, while recall measures the proportion of truly included studies that were correctly identified by the model. These metrics jointly assess both correctness and completeness of model screening decisions.
B.3 Evidence summarization
Definition & Motivation.
This task simulates the final synthesis step of a systematic review, in which the model must draw a conclusion from the included studies’ findings. Given a self-contained multiple-choice question based on a Cochrane review’s objectives, the model selects the answer option that aligns with the review’s main conclusion. An example instance is shown in Supplementary Table 9.
Input & Output.
The input to the evidence summarization task consists of a self-contained multiple-choice question generated from a Cochrane review abstract (including its answer options A, B, C, …) together with the titles and full structured abstracts of all studies included in that review (ranging from 1 to 20 PMIDs). The model’s output is a single label (A, B, C, …) corresponding to the correct conclusion. Successfully performing this task requires the model to comprehend scientific text, aggregate evidence across multiple documents, and map study findings to the correct conclusion.
Dataset Construction.
First, Cochrane review abstracts are processed by GPT-4o using a prompt that instructs the model to generates a self-contained multiple-choice question based on the review objectives, to produce answer options covering plausible outcomes, and to output only valid JSON with the keys question, options, and answer. The prompt also forbids referencing the original review by phrases such as “based on” to ensure questions are fully self-contained, and the model is run at temperature 0 to guarantee deterministic outputs. Next, for each generated question, we retrieve the included-study PMIDs from the review metadata and load the corresponding titles and structured abstracts from our PubMed cache, discarding any entries with zero or more than twenty included studies. Finally, the remaining $N$ examples are sorted by numeric PMID (serving as a proxy for publication date); the most recent 1 000 examples form the test set, the preceding 500 form the validation set, and the remainder form the training set.
Evaluation Metrics.
We report accuracy and macro-F1 to account for class imbalance across answer choices.
Table 9: Example instance of Evidence Summarization
Source PubMed ID: 10796093 Question: What were the findings regarding the effectiveness of bed rest in hospital for women with suspected impaired fetal growth compared to ambulatory management? Please answer this question based on the information provided below: { title and abstract of included studies } Options: A. Bed rest in hospital significantly improved fetal growth parameters and neonatal outcomes. B. Bed rest in hospital significantly worsened fetal growth parameters and neonatal outcomes. C. There were no significant differences in fetal growth parameters and neonatal outcomes between bed rest in hospital and ambulatory management. D. Bed rest in hospital was found to be harmful to both the mother and the fetus. Answer: C
Appendix C Benchmark building: Trial design tasks
C.1 Arm design
Definition & Motivation.
This task assesses whether the model can identify the correct treatment arms and intervention details for a given clinical trial. Based on the trial’s title and brief summary, the model must choose among four candidate descriptions the one that accurately represents the trial’s arm labels, arm types, and assigned interventions. An example instance is shown in Supplementary Table 10.
Input & Output.
The model is provided the official trial title and summary, followed by four arm descriptions that list the group labels, intervention types, and associated treatments. It must return the label corresponding to the correct description.
Dataset Construction.
We start with sets of up to four trial identifiers drawn from trials cited together in the same Cochrane review. For each identifier, a script extracts the arm groups and individual intervention details from our trial metadata. If fewer than four candidates are available, additional trials are added at random (with a fixed seed for reproducibility). To create challenging distractors, incorrect descriptions are drawn from the same arm and intervention information of other trials cited by that review. All items are then sorted by the numeric portion of the trial identifier and partitioned so that the latest 1,000 trials form the test set, the next 500 form validation, and the rest form training.
Evaluation Metrics.
We report accuracy and macro-F1 to account for class imbalance across answer choices.
Table 10: Example instance of Trial Arm Selection
Source NCTID: NCT00000150 Question: Please select the arm or intervention descriptions that belong to the clinical trial below. Title: Submacular Surgery Trials (SST) Brief Summary: To determine whether surgical removal of subfoveal choroidal neovascularization (CNV) and associated hemorrhage in patients with age-related macular degeneration (AMD), the ocular histoplasmosis syndrome (OHS), or idiopathic CNV stabilizes or improves vision more often than observation. To determine how surgical removal compared to observation of subfoveal CNV due to AMD, OHS, or idiopathic causes changes the patient’s perception of health- and vision-related “quality of life,” as measured by telephone interview using the Medical Outcomes Survey Short Form-36 (MOS SF-36), the Hospital Anxiety and Depression Scale, and the National Eye Institute Visual Function Questionnaire (NEI VFQ-25). To determine whether randomized trials of surgery are warranted for patients with subfoveal CNV associated with AMD not suitable for laser treatment. Options: A. Intervention Type: DRUG; Name: rhuFab V2 (ranibizumab); Assigned to Arm(s): N/A B. Arm Label: bevacizumab; Type: EXPERIMENTAL; Interventions: Drug: Bevacizumab
Arm Label: control; Type: NO_INTERVENTION; Interventions: N/A
Intervention Type: DRUG; Name: Bevacizumab; Assigned to Arm(s): bevacizumab
C. Intervention Type: PROCEDURE; Name: Subfoveal Choroidal Neovascularization Removal; Assigned to Arm(s): N/A D. Intervention Type: DRUG; Name: rhuFab V2 (ranibizumab); Assigned to Arm(s): N/A
Answer: C
C.2 Eligibility criteria design
Definition & Motivation.
This task measures the model’s ability to match a trial’s inclusion and exclusion criteria. Given a trial’s title and summary, the model must select, from four text options, the one that accurately lists the trial’s eligibility rules. An example instance is shown in Supplementary Table 11.
Input & Output.
The model receives the trial title and summary along with four candidate eligibility criterion sets. It must output the correct option label.
Dataset Construction.
For eligibility criteria, dataset construction follows the same approach as in Arm Design: we form candidate sets from trials cited by the same Cochrane review, extract each trial’s eligibility rules from our metadata, pad with random trials when needed, and draw distractors from other cited trials to ensure relevance. Examples are then sorted by the numeric part of the trial identifier and split into test (latest 1,000), validation (next 500), and training (remainder).
Evaluation Metrics.
We report Accuracy and macro-F1 to account for class imbalance across answer choices.
Table 11: Example instance of Trial Eligibility Criteria Selection
Source NCTID: NCT00004563 Question: Please select the eligibility criteria that correspond to the clinical trial below. Title: Cyclophosphamide Versus Placebo in Scleroderma Lung Study Brief Summary: To evaluate the efficacy and safety of cyclophosphamide versus placebo for the prevention and progression of symptomatic pulmonary disease in patients with systemic sclerosis. Options: A. Inclusion: systemic scleroderma with active alveolitis on BAL fluid (neutrophilia >=3% or eosinophilia >=2%) on CT; symptom onset (excluding Raynaud’s) within 7 years; FVC 45–85% predicted; Mahler Dyspnea Index grade 2; Exclusion: DLCO <30% predicted; smoking in past 6 months; pulmonary abnormalities; pulmonary hypertension requiring therapy; prednisone >10 mg/day or prior cyclophosphamide >4 weeks or >=2 IV doses; recent disease-modifying meds; … B. Inclusion: idiopathic pulmonary fibrosis per ATS 2011, duration 6–48 months; age 40–80; %FVC >=50% <=90%; %DLCO >=30% <=90%; Exclusion: FEV1/FVC <0.8 post-bronchodilator; expected lung transplant within 1 year or on waiting list; known ILD cause; asthma/COPD history; active infection; ongoing IPF treatments; unstable cardiac/pulmonary disease in past 6 months; … C. Inclusion: IPF diagnosis; age 40–80; FVC >=50% predicted; DLCO >=35% predicted; FVC or DLCO <=90% predicted; no improvement in past year; walk 150 m in 6 min with sat >=83% on <=6 L/min O2; Exclusion: unable PFT; obstructive lung disease or airway hyper-responsiveness; expected lung transplant within 72 weeks; active infection; liver disease; cancer or other condition likely death within 2 years; diabetes; pregnancy/lactation; substance abuse; long QT syndrome; other IPF treatment; inability to take study medication; withdrawal from IPF trials; … D. Inclusion: IPF diagnosis; age 40–80; FVC >=50% predicted; DLCO >=35% predicted; FVC or DLCO <=90% predicted; no improvement in past year; walk 150 m in 6 min with sat >=83% on <=6 L/min O2; Exclusion: unable PFT; obstructive lung disease or airway hyper-responsiveness; expected lung transplant within 72 weeks; active infection; liver disease; cancer or other condition likely death within 2 years; diabetes; pregnancy/lactation; substance abuse; long QT syndrome; other IPF treatment; inability to take study medication; withdrawal from IPF trials; … Answer: A
C.3 Endpoint design
Definition & Motivation.
This task evaluates whether the model can identify the primary outcome measures of a clinical trial. Based on the trial title and summary, the model must pick the option that lists the correct primary endpoints. An example instance is shown in Supplementary Table 12.
Input & Output.
The input consists of the trial title and summary plus four outcome descriptions that concatenate the primary measures and any associated timeframes. The model must choose the correct label.
Dataset Construction.
For primary endpoints, dataset construction also follows the Arm Design template: trials cited together provide candidate sets, we extract each trial’s primary outcomes from our metadata, pad with random examples if needed, and sample distractor options from other cited trials. The resulting items are ordered by the numeric portion of the trial identifier and divided into test (latest 1,000), validation (next 500), and training splits.
Evaluation Metrics.
We report Accuracy and macro-F1 to account for class imbalance across answer choices.
Table 12: Example instance of Trial Outcome Measure Selection
Source NCTID: NCT00000170 Question: Please select the outcome measures that correspond to the clinical trial below. Title: Amblyopia Treatment Study: Occlusion Versus Pharmacologic Therapy for Moderate Amblyopia Brief Summary: To determine whether success rate with atropine vs. patching therapy in <7 y olds with strabismic or anisometropic amblyopia is equivalent; To develop precise estimates of treatment success rates; To identify factors associated with successful treatment; To collect detailed amblyopia course data; Extended follow-up: Primary—long-term visual acuity at age 10 and 15; Secondary—compare long-term acuity at age 10 and 15 between patching+care vs. atropine+care. Options: A. Visual acuity of amblyopic eye by HOTV (<7 y) or ETDRS (>=7 y); measurement at 2 wk (strabismic/anisometropic/combined) and 4 wk (deprivation); … B. Maximal improvement in visual acuity; … C. Change in visual acuity from baseline to week 6 post-randomisation between arms; 6 wk; … D. Visual acuity in amblyopic eye at 6 months; … Answer: D
C.4 Sample size estimation
Definition & Motivation.
This task evaluates the model’s quantitative reasoning by asking it to estimate the number of participants required for a clinical trial based on its design and stated statistical assumptions. Success requires understanding trial objectives, translating textual assumptions (effect size, significance level, power, dropout) into a numerical enrollment target, and applying standard sample-size formulas. An example instance is shown in Supplementary Table 13.
Input & Output.
As input, the model receives: a narrative description of the trial design—official title, conditions, arm groups, interventions and primary outcomes, as well as a concise summary of the statistical assumptions extracted from the protocol text (treatment effect, alpha, power, expected dropout). The model must output a single integer corresponding to the estimated sample size.
Dataset Construction & Splits.
We linked 5,737 PubMed protocols to their NCTIDs and identified 4,315 full-text sections titled “sample size” together with 5,678 CT.gov records containing enrollment counts. Matching the reported CT.gov enrollment with the corresponding paper section yielded 2 143 examples where the protocol paragraph explicitly mentions the official enrollment number. For each example, we used a deterministic prompt to summarize the paragraph into (1) effect-size assumptions, (2) statistical parameters (significance level, power, dropout rate), and (3) the true enrollment. We then generated a question by combining trial metadata with these summarized assumptions. Finally, all examples were sorted by the numeric part of the NCTID; the most recent 1,000 form the test set, the next 500 the validation set, and the remainder the training set.
Evaluation Metrics.
We report (1) the mean absolute error between predicted and actual enrollments, and (2) the proportion of trials for which the model’s estimate lies within ±20 % of the true sample size.
Table 13: Sample Size Estimation Question
Source NCTID: NCT06330298 Question: Here is the design of a clinical trial: Official Title: Improving Social Cognition and Social Behaviour in Various Brain Disorders. Included conditions: Stroke; Multiple Sclerosis; Brain Tumor. Study Arm groups: Experimental condition: Receives T-ScEmo Treatment (type EXPERIMENTAL; receives T-ScEmo between T0 and T1; interventionNames: [“Behavioral: Treatment social cognition and emotion regulation (T-ScEmo)”]); Waiting list group: Will be on waiting list instead of treatment (type NO INTERVENTION; duration T0–T1). Interventions: Type BEHAVIORAL; name Treatment social cognition and emotion regulation (T-ScEmo); description T-ScEmo is a multifaceted protocol to improve social cognition, regulation of social behaviour and participation (Westerhof et al., 2017, 2019); 15×1 h live sessions + 5 online practice sessions; involves significant other and homework; armGroupLabels: [Experimental condition: Receives T-ScEmo Treatment]. Primary Outcomes: Change in social behaviour examined by proxy; difference T0→T2 on the Dysexecutive Questionnaire Social scales proxy version (DEX-Socproxy; Spikman et al., 2013; Westerhof-Evers, 2017); time frame: through completion (8–10 months). Please estimate the sample size based on: significance level ($\alpha=0.05$), power ($0.801-\beta=$ 0.80), and the group×$\times$×time interaction in a repeated measures analysis. Answer: 84 Explanation: The study of Westerhof-Evers et al. [9] found that the DEX-Socproxy had an effect size of Cohen’s d = 0.2 (Cohen’s f = 0.10) for the comparison between T0 and T2. From these data, a small to medium effect size of Cohen’s f = 0.15 was estimated for the current study. Therefore, the sample size is based on a power calculation for the main outcome measure (DEX-Socproxy) regarding the condition $\times$ time interaction (T0, T2) using repeated measures analysis with effect size Cohen’s f = 0.15, $\alpha=0.05$, and power = 0.80 [29], resulting in a sample size of $n=62$. However, to ensure balanced distribution across patient groups, a total sample size of $n=84$ will be used (42 patients per condition).
C.5 Trial completion assessment
Definition & Motivation.
This task simulates a real-world feasibility analysis during clinical trial planning, where the model must assess whether a trial is likely to reach completion or be terminated early based on its protocol design. In addition, if a trial is predicted to terminate, the model is asked to identify the most likely reason for termination from a predefined set of categories (e.g., enrollment issues, safety concerns, lack of efficacy, operational problems). This reflects practical needs in the early design stage when sponsors assess trial viability. An example instance is shown in Supplementary Table 14.
Input & Output.
Each example consists of structured metadata and protocol text for a registered interventional study, including trial phase, number of arms, masking, intervention model, primary purpose, arm descriptions, and detailed eligibility criteria. The output is a binary label indicating whether the trial was completed or terminated. For trials labeled as terminated, the model must also predict the most likely termination reason from one of several standardized categories.
Dataset Construction.
We construct the dataset using completed and terminated interventional trials from ClinicalTrials.gov with known outcomes. Structured fields are extracted from the studies, disposition, eligibility_criteria, and arms tables in our trial database. Only trials with definitive outcome labels (i.e., Completed or Terminated) and non-missing eligibility criteria are retained. Termination reasons are normalized into broad categories using a curated taxonomy (e.g., enrollment issues, lack of efficacy, safety, sponsor decision). To avoid label leakage, we exclude any outcomes or termination reasons from the input text fields.
Evaluation Metrics.
We evaluate trial completion prediction using accuracy. For terminated trials, termination reason classification is evaluated using macro-F1 to balance across categories with different frequencies.
Table 14: Example instance of Trial Completion Assessment
Source NCTID: NCT02823470 Study Title: A Pilot Study to Evaluate the Use of Smart Adherence Technology to Measure Lumacaftor/Ivacaftor Adherence in CF Subjects Homozygous for the F508del CFTR Mutation Study Design: Allocation: Randomized Intervention Model: Parallel Masking: None Primary Purpose: Treatment Phase: Phase 4 Number of Arms: 2 Arms: Arm A (Experimental): Activated smart device alerts and feedback
LUM/IVA: LUM 400 mg q12h / IVA 250 mg q12h through Week 48.
Arm B (Experimental): De-activated smart device alerts/feedback features
LUM/IVA: LUM 400 mg q12h / IVA 250 mg q12h through Week 48.
Inclusion Criteria: CF diagnosis with F508del-CFTR homozygosity, FEV1 $\geq$ 40% predicted, consented, willing to comply. … [truncated for brevity] Exclusion Criteria: Hepatic impairment, mechanical ventilation, pregnancy, recent investigational drug use, prior/current use of commercial LUM/IVA. … [truncated for brevity] Outcome: Terminated Reason for Termination: Enrollment issues — study terminated early due to slow recruitment of patients.

### v2

@@
-TrialPanorama: Database and Benchmark for Systematic Review and Design of Clinical Trials
+Developing Large Language Models for Clinical Research Using One Million Clinical Trials
 Abstract
-Developing artificial intelligence (AI) for vertical domains requires a solid data foundation for both training and evaluation.
-In this work, we introduce TrialPanorama, a large-scale, structured database comprising 1,657,476 clinical trial records aggregated from 15 global sources.
-The database captures key aspects of trial design and execution, including trial setups, interventions, conditions, biomarkers, and outcomes, and links them to standard biomedical ontologies such as DrugBank and MedDRA.
-This structured and ontology-grounded design enables TrialPanorama to serve as a unified, extensible resource for a wide range of clinical trial tasks, including trial planning, design, and summarization.
-To demonstrate its utility, we derive a suite of benchmark tasks directly from the TrialPanorama database.
-The benchmark spans eight tasks across two categories: three for systematic review (study search, study screening, and evidence summarization) and five for trial design (arm design, eligibility criteria, endpoint selection, sample size estimation, and trial completion assessment).
-The experiments using five state-of-the-art large language models (LLMs) show that while general-purpose LLMs exhibit some zero-shot capability, their performance is still inadequate for high-stakes clinical trial workflows.
-We release TrialPanorama database and the benchmark to facilitate further research on AI for clinical trials.
-1 Introduction
-Developing new drugs is a time-consuming and costly endeavor, often requiring up to a decade and exceeding \$2 billion in expenses [1].
-A substantial portion of this burden stems from the human study phase, during which clinical trials are conducted to assess safety and efficacy across phases I through IV.
-In addition, clinicians, researchers, and trial designers routinely search and review clinical trial records and results.
-These efforts support critical objectives such as designing new trials, evaluating treatments, and informing clinical decision-making in evidence-based medicine [2, 3].
-Clinical trial records are dispersed across global registries, including ClinicalTrials.gov [4], the International Clinical Trials Registry Platform (ICTRP) [5], the EU Clinical Trials Register [6], and others [7, 8].
-Additionally, many trials are reported in the literature, particularly on PubMed.
-Prior research has focused on curating subsets of this data for specific objectives, such as clinical trial results [9, 10], drug and condition annotations [11, 12], trial-publication linkages [13], and structured outcomes [14].
-A unified view that integrates these diverse aspects of clinical trial data can substantially improve evidence synthesis in systematic reviews and enable more informed, data-driven approaches to designing future clinical trials [15, 16, 17].
-In this work, we build TrialPanorama, a clinical trial database that integrates multiple data sources and standardizes trial-level design and result attributes, aiming to provide a great resource for clinical trial design and review as well as developing AI models for these tasks.
-Artificial intelligence (AI), particularly with the emergence of large language models (LLMs), has garnered significant attention in clinical trial applications.
-Recent advances have explored tasks including predicting clinical trial outcomes [18, 19, 20], matching patients to eligible trials [21, 22], extracting cohort real-world data [23, 24], writing clinical trial documents [25, 26], and accelerating systematic review of the literature of clinical studies [27, 28].
-However, these developments are often limited by the lack of standardized, comprehensive clinical trial data suitable for AI development.
-TrialPanorama bridges this gap by (1) offering curated benchmark data to support the development and evaluation of AI methods for trial applications, and (2) providing source-grounded evidence to enhance AI for trial applications through retrieval-augmented generation (RAG) [29].
-Table 1: Comparison of TrialPanorama and prior clinical trial datasets and benchmarks.
-Columns: Size = dataset size (# trials), Src = data source (CTG = ClinicalTrials.gov), Des = structured trial design data (e.g., arms, eligibility), Res = structured trial results (e.g., adverse events), Out = trial outcomes, Rel = trial-study/publication/review relations, Ont = linked to medical ontologies (e.g., DrugBank, MedDRA), Bench-D = trial design benchmark tasks, Bench-SLR = systematic literature review benchmark tasks.
-Dataset / Benchmark Size Src Des Res Out Rel Ont Bench-D Bench-SLR
-TrialPanorama (Ours) 1.6M Multi(15 sources)
-TOP [19] 10K CTG
-CTKG [11] 8K CTG
-CTGResults [9] 56K CTG
-CTO [14] 125K CTG
-TrialBench [30] 400K CTG
-CARE [10] 700 pubs PubMed
-PubMedKG [31] 480K PubMed,CTG
-TrialReviewBench [28] 100 reviews PubMed,CTG
-LEADSInstruct [27] 21K reviews PubMed,CTG
-In this work, we introduce TrialPanorama, the largest-to-date structured clinical trial database, integrating over two million records from 16 global sources, including clinical trial registries, biomedical literature, and systematic reviews.
-As illustrated in Figure 1a, the database captures essential elements of clinical trials, such as study designs, interventions, conditions, outcomes, and biomarkers, and standardizes them through mappings to biomedical ontologies like DrugBank and MedDRA.
-Building on this data foundation, we further demonstrate how TrialPanorama enables tasks that support AI development for trial design and systematic review.
-These tasks span both upstream evidence synthesis (e.g., study search, screening, summarization) and downstream trial planning (e.g., arm design, eligibility criteria, endpoint selection, sample size estimation, and trial completion prediction), as illustrated in Figures 1b and  1c.
-Compared to existing trial datasets and benchmarks (Table 1), TrialPanorama offers broader coverage, richer structure, and stronger alignment with real-world AI use cases in clinical research.
-We release the database and benchmark to facilitate future research in this space.
-Figure 1: Overview of the TrialPanorama dataset and benchmark suite.
-(a) Schema of the structured clinical trial database, showing linked tables for trial protocols, interventions, outcomes, and metadata.
-(b) Summary of the systematic review benchmark tasks.
-(c) Summary of the clinical trial design benchmark tasks.
-2 TrialPanorama: Database
-As illustrated in Figure 1a, the TrialPanorama database comprises ten main tables: studies, conditions, drugs, endpoints, relations, biomarkers, outcomes, adverse events, results, and relations.
-These tables are organized into four conceptual clusters: trial attributes, trial protocols, trial results, and study-level links.
-The trial metadata cluster captures core metadata such as study title, brief summary, sponsor, start year, and recruitment status.
-The trial protocol cluster describes the setup and design of the study, including tested drugs, targeted conditions, and patient group allocations.
-The trial results cluster encompasses reported findings, including the trial outcomes, adverse events, and efficacy results.
-The study-level link cluster encodes relationships across studies, such as mappings between registry records and corresponding publications, or reviews aggregating multiple studies on a common clinical topic.
-In the following subsections, we describe the content of each table in detail.
-Data acquisition and processing procedures are provided in Appendix A.
-2.1 Trial metadata
-The studies table serves as the central registry in TrialPanorama, containing 1,657,476 clinical trial records from 15 diverse sources, including ClinicalTrials.gov, PubMed, CHICTR, etc.
-The distribution of the sources is illustrated in Figure 2a.
-Each row, uniquely indexed by a study_id, captures metadata such as trial type (INTERVENTIONAL, OBSERVATIONAL), sponsor (e.g., university, hospital, pharmaceutical), and temporal attributes including start_year, enabling longitudinal analysis.
-The table also includes regulatory fields like phase (e.g., PHASE1PHASE4) and recruitment status (e.g., RECRUITING, COMPLETED, SUSPENDED), providing a standardized overview of the trial’s attributes.
-2.2 Trial protocols
-The trial protocol cluster in TrialPanorama captures structured information describing the design and setup of clinical studies, including target populations, interventions, outcomes, and biological markers.
-This cluster comprises five key tables: drugs, conditions, biomarkers, dispositions, and endpoints.
-Figure 2: Overview of the database statistics.
-a, Distribution of clinical trials across major data sources.
-b, Frequency distribution of the top 100 drugs mentioned in the database, with the most common drug classes annotated.
-c, Frequency distribution of the top 100 medical conditions, annotated by condition categories.
-Drugs.
-The drugs table records the therapeutic interventions evaluated across trials, comprising 1,498,985 rows and 15 columns.
-Each entry is linked to a study_id, with over 616,000 unique studies represented.
-The distribution of top frequent drugs is shown in Figure 2b.
-The drug_name field captures 404,175 unique names, including both generic labels (e.g., “Placebo”) and brand/composite entries.
-To enable structured pharmacological analysis, drugs are linked to DrugBank identifiers (drugbank_id) and normalized RxNorm terms (rx_normalized_name).
-Mechanism of action (MoA) information is encoded via the drug_moa_id field, which maps each drug to its biological targets, such as enzymes, receptors, or transporters, and their interaction types (e.g., inhibitor, agonist).
-The definitions and details of each MoA are stored in a separate drug_moa table, where drug_moa_id serves as the key linking back to the drugs table.
-Conditions.
-The conditions table catalogs the disease or physiological focus of each study, containing 10,133,060 entries from over 1,074,458 studies.
-Figure 2c shows the most frequent conditions.
-The condition_name field includes 5,236 unique MeSH-derived terms representing medical conditions, while condition_mesh_id ensures mapping to standardized biomedical ontologies.
-These annotations facilitate the retrieval and grouping of studies by clinical focus (e.g., cardiovascular or neurological diseases).
-Biomarkers.
-The biomarkers table records therapeutic biomolecular and clinical markers used to stratify, monitor, or predict treatment response.
-It contains 477,717 rows across 339,894 unique studies.
-Each biomarker is annotated by its biomarker_type (e.g., predictive, surrogate endpoint, safety), biomarker_name, and associated gene markers (biomarker_genes).
-This table supports downstream applications such as biomarker-driven subgroup analysis or treatment personalization.
-Dispositions.
-The dispositions table describes how participants are grouped and assigned to interventions within each study.
-With over 4.5 million entries covering 1.47 million unique studies, it specifies the intervention_type (e.g., DRUG, DEVICE, BEHAVIORAL), group_type (e.g., EXPERIMENTAL, COMPARATOR, PLACEBO), and intervention_name.
-This table is critical for reconstructing arm-level trial structure and modeling allocation strategies across diverse study designs.
-Endpoints.
-The endpoints table defines the primary clinical outcomes of interest across 2,621,226 rows and 1.38 million trials.
-Each entry includes a primary_endpoint label (e.g., Overall Survival, Quality of Life), categorized into high-level primary_endpoint_domain (e.g., mortality, physiological) and fine-grained primary_endpoint_subdomain (e.g., cardiovascular outcomes, emotional well-being).
-2.3 Trial results
-The trial results cluster captures the observed clinical outcomes reported at the end of each study, including both efficacy and safety measures.
-It comprises three main tables: outcomes, results, and adverse_events, each offering distinct perspectives on the impact of tested interventions.
-Outcomes.
-The outcomes table contains 532,525 entries, each uniquely mapped to a study.
-It records overall trial status (overall_status), such as completed, recruiting, or terminated, and provides a high-level assessment of outcome success through the outcome_type field (e.g., completed, positive outcome/primary endpoint(s) met, unknown, or terminated, other/unclear).
-These annotations facilitate downstream analyses such as success rate estimation and status tracking across therapeutic areas and time.
-Results.
-The results table comprises over 5.7 million entries covering more than 1.1 million studies.
-It records arm-level trial outcomes, including structured descriptions of study populations (population), tested interventions (interventions), observed clinical outcomes (outcomes), and treatment group type (group_type).
-The granularity of eligibility criteria and baseline characteristics captured in the population field supports fine-grained subgroup analysis, while the group_type distinction (e.g., EXPERIMENTAL, COMPARATOR) enables comparative outcome interpretation across arms.
-Adverse Events.
-The adverse_events table documents reported safety events across 6,145,012 rows and over 530,000 unique studies.
-It captures adverse_event_name (e.g., Headache, Nausea, Diarrhoea) along with corresponding MedDRA-coded identifiers (meddra_id), enabling standardized aggregation and harmonization of safety profiles across studies.
-These data support pharmacovigilance and risk-benefit assessments for individual therapies or therapeutic classes.
-2.4 Study links
-The relations table captures cross-study relationships that connect clinical trial records from different sources or define higher-order linkages such as reviews or citation networks.
-This table contains 940,582 relation instances and supports integration, traceability, and synthesis of trial-level evidence.
-Each row represents a directed edge between a head_id and a tail_id, with associated head_type and tail_type identifying the origin of the studies (e.g., PubMed, ClinicalTrials.gov, or regional registries).
-This structure allows the database to represent connections between registry entries and publications describing the same study, or group related records across different national databases.
-The semantic type of each link is encoded in the relation_type field, which includes labels such as linked_to (e.g., registry-publication pair), cite (e.g., PubMed-to-PubMed citation), include (e.g., a review including a trial), and exclude (e.g., a review explicitly excluding a study).
-These relationships enable downstream use cases such as study clustering, citation graph construction, and synthesis of trial networks for systematic reviews.
-Figure 3: Manual data quality assessment across key database tables. The y-axis shows the proportion of valid (i.e., correctly extracted) records. Sample sizes (n) are shown below each bar.
-2.5 Data quality assessment
-To assess the quality of TrialPanorama, we employed seven annotators to manually verify approximately 1,500 records sampled across all major tables, stratified by data sources.
-Each sampled record was evaluated against the original source to determine its validity.
-As summarized in Figure 3, the database exhibits high overall accuracy, with most tables showing a validity rate above 90%.
-Notably, errors observed in some tables, particularly conditions and biomarkers, can be attributed to challenges in text-based pattern matching and the use of LLMs for information extraction.
-Despite these limitations, the results highlight the feasibility and promise of using LLMs to automatically structure clinical trial data from heterogeneous and unstructured sources.
-3 TrialPanorama: Benchmark
-TrialPanorama serves as a valuable resource for a wide range of clinical trial applications by offering a global view of trial protocols, outcomes, and systematic reviews.
-Building on this comprehensive foundation, we develop a suite of representative benchmark tasks spanning two key application areas: three tasks focused on systematic evidence synthesis and five tasks supporting data-driven clinical trial design.
-These benchmarks are illustrated in Figures 1b and 1c.
-3.1 Systematic review tasks
-The PRISMA statement [32] outlines standard procedures for conducting systematic reviews and meta-analyses in medicine.
-Among these, three critical steps are: (1) study search: identifying candidate studies from literature databases; (2) study screening: selecting eligible studies based on inclusion and exclusion criteria; and (3) evidence synthesis: aggregating findings from included studies to answer a targeted clinical question, such as the efficacy of a therapy for a specific condition.
-Based on TrialPanorama, we construct benchmark tasks corresponding to each of these key steps: study search, study screening, and evidence summarization.
-By leveraging the relations table, we are able to identify linkages between systematic review articles and the clinical trials they include or exclude, enabling the construction of high-quality, labeled benchmarks for these tasks.
-We hold the most recent studies and reviews out as the test set.
-The details of benchmark building, setups, and example samples are in Appendix B.
-Study Search.
-The study search task evaluates an AI model’s ability to retrieve relevant clinical trials from a large corpus, given the structured context of a systematic review.
-It simulates the first PRISMA step: identifying candidate studies based on a review’s topic.
-Each benchmark instance includes a query review defined by its background, objectives, and selection criteria, and models must retrieve trials that were included in the final review.
-Ground-truth labels are derived from the relations table in TrialPanorama.
-The benchmark includes 8,004 training reviews linked to 175,577 studies and 100 test reviews with 4,324 relevant studies.
-Retrieval performance is measured using Recall@K.
-Study Screening.
-The study screening task evaluates an AI model’s ability to determine whether candidate studies meet the inclusion criteria of a systematic review, simulating the second PRISMA step of study selection.
-Each benchmark instance includes a structured review query (background, objectives, and criteria) and 10 candidate studies described by both unstructured text (title and abstract) and structured metadata from TrialPanorama (e.g., phase, intervention, conditions).
-Ground-truth labels are derived from explicit inclusion/exclusion links in the relations table.
-Notably, the excluded candidates represent hard negatives, as they were manually excluded by the original review authors and recorded in the review’s study list.
-The benchmark includes 7,931 training reviews with over 175,000 included and 229,000 excluded studies, and 100 test reviews with 491 included and 488 excluded studies.
-Performance is evaluated using accuracy, precision, and recall.
-Evidence Summarization.
-The evidence summarization task evaluates an AI model’s ability to answer clinical research questions by synthesizing findings from relevant studies, emulating the last PRISMA step: evidence summarization.
-Each instance includes (1) the review’s structured setup (background, objectives, selection criteria), (2) a set of study abstracts containing relevant or irrelevant evidence, and (3) a multiple-choice question summarizing possible conclusions.
-Input data and gold labels are constructed using the relations and results tables in TrialPanorama, where study-review links define the context and structured trial findings inform the correct answer.
-Answer choices reflect typical clinical conclusions (e.g., “limited evidence” or “effective with adverse effects”).
-The benchmark contains 4,636 reviews and 35,512 studies in the training set, and 1,000 reviews with 8,974 studies in the test set.
-Model performance is assessed using accuracy and macro-F1 score.
-3.2 Clinical trial design tasks
-Designing a clinical trial involves multiple interdependent components that collectively determine its scientific validity and operational feasibility.
-As illustrated in Figure 1c, key design tasks include: (1) sample size estimation, which determines the number of participants needed to ensure sufficient statistical power to detect treatment effects; (2) eligibility criteria design, which defines the characteristics of the target population eligible for inclusion; (3) arm design, which specifies the allocation of participants into treatment and control groups with defined interventions; and (4) endpoint design, which identifies the clinical measures used to assess efficacy and safety.
-In addition, trial completion assessment evaluates whether a trial is likely to reach completion or be terminated early due to design flaws, recruitment challenges, or safety concerns, thus providing a holistic perspective on the robustness of the trial design.
-We hold the most recent studies out as the test set.
-The details of benchmark building, setups, and example samples are in Appendix C.
-Arm, Eligibility Criteria, and Endpoint Design.
-The trial design tasks assess a model’s ability to identify optimal design components for a given trial protocol.
-Each task presents a structured input, consisting of the trial’s title and brief summary, and asks the model to select the correct choice from a set of candidate design elements, where distractors are sampled from similar trials to ensure difficulty.
-In the arm design task, the goal is to select the correct configuration of intervention groups; in eligibility criteria design, the correct inclusion and exclusion conditions; and in endpoint design, the appropriate outcome measures.
-Candidate options are drawn from structured records in TrialPanorama, including data from the studies, dispositions, conditions, and endpoints tables.
-For all three tasks, performance is evaluated using classification accuracy.
-The benchmark includes 27,321 training studies and 1,000 test studies for each of the three tasks.
-Sample Size Estimation.
-The sample size estimation task evaluates a model’s ability to infer the required number of participants for a clinical trial, based on its structured protocol, including target conditions, interventions, and primary outcomes, under a fixed drop-out assumption.
-Each instance draws data from the studies, dispositions, and endpoints tables in TrialPanorama, paired with gold-standard target values extracted from protocols or publications.
-The benchmark includes 1,143 training examples and 1,000 test studies.
-Model performance is assessed using two metrics: accuracy (percentage of estimation within ±20% of ground truth) and mean absolute error (MAE).
-This task requires both clinical reasoning and quantitative estimation skills grounded in trial design logic.
-Trial Completion Assessment.
-This task evaluates a model’s ability to predict whether a clinical trial will reach completion or terminate early based on its protocol, and, if terminated, to identify the most likely reason (e.g., efficacy, safety, enrollment failure).
-Each instance includes structured trial metadata, such as study phase, arm count, masking, allocation, and eligibility criteria, drawn from the studies, dispositions, and results tables in TrialPanorama.
-The benchmark comprises 42,195 training studies and 1,000 test studies.
-Model performance is evaluated by balanced accuracy on (1) the binary completion prediction and (2) multiclass termination reason classification.
-This task reflects real-world feasibility assessments during trial planning.
-4 Experiment
-We selected a diverse set of state-of-the-art proprietary and open-source LLMs for evaluation.
-These include GPT-4o and its lightweight variant GPT-4o-mini, as well as the recent reasoning-focused model O3-mini, all accessed via the Azure OpenAI Service.
-In addition, we include the LLaMA-3.3-70B-Instruct and LLaMA-3.1-8B-Instruct models, representing large and moderate-scale open-source LLMs, accessed through the Azure AI Inference Service.
-This selection allows us to compare models across different scales, architectures, and reasoning capabilities.
-Table 2: Benchmark experiment results for systematic review tasks.
-Tasks Study search Study screening Evidence summarization
-Models Recall@100 Recall@500 Accuracy Precision Recall Accuracy Macro-F1
-Llama-8B 11.6 16.0 55.3 46.8 58.2 70.4 71.1
-Llama-70B 25.1 33.3 60.7 59.1 70.5 78.8 76.7
-GPT-4o-mini 17.3 22.7 60.2 60.4 60.5 74.1 73.1
-GPT-4o 22.9 27.3 66.1 64.4 72.3 72.8 65.2
-O3-mini 27.6 37.8 71.6 71.9 71.3 72.7 53.8
-4.1 Systematic review tasks
-Table 2 presents the performance of various LLMs across three core systematic review tasks: study search, study screening, and evidence summarization.
-The results highlight both general limitations and comparative strengths across models.
-First, performance on the study search task is uniformly low, with Recall@100 ranging from 11.6 (LLaMA-8B) to 27.6 (O3-mini).
-This reflects a key weakness of general-purpose LLMs: they struggle to generate precise and discriminative search queries that can retrieve relevant trials from a large corpus.
-Notably, O3-mini outperforms both proprietary (GPT-4o, GPT-4o-mini) and open-source (LLaMA-70B, LLaMA-8B) models by a significant margin, suggesting its reasoning-focused training may better align with the query formulation challenge.
-Second, study screening performance is moderate, with accuracy scores ranging from 55.3 (LLaMA-8B) to 71.6 (O3-mini).
-The task is intrinsically difficult, as the negative examples in the test set are hard negatives.
-These studies are closely related to the review but were explicitly excluded by the original systematic review authors.
-Despite this challenge, O3-mini demonstrates the strongest overall performance in accuracy, precision, and recall, followed by GPT-4o.
-LLaMA-70B shows reasonable performance, while LLaMA-8B significantly underperforms.
-Third, in evidence summarization, model performance varies more widely.
-LLaMA-70B achieves the highest accuracy (78.8) and Macro-F1 (76.7), outperforming GPT-4o and GPT-4o-mini, which show moderately strong but less consistent results.
-Surprisingly, O3-mini, which performs best on study search and screening, shows notably weak performance in summarization (Macro-F1 = 53.8), despite being positioned as a reasoning-oriented model.
-This result is somewhat counterintuitive, as reasoning-capable models are typically expected to excel at inference tasks that involve aggregating evidence from multiple sources.
-Table 3: Benchmark experiment results for clinical trial design tasks.
-“Acc”: short for accuracy; “BAcc-out”: short for balanced accuracy of trial outcome assessment results; “BAcc-rt”: short for balanced accuracy of trial termination rationalization results; “MAE $\downarrow$”: mean absolute error that is the smaller the better.
-Tasks Arm design Eligiblity criteria design Endpoint design Sample size estimation Trial completion
-Models Acc Macro-F1 Acc Macro-F1 Acc Macro-F1 Acc MAE $\downarrow$ BAcc-out BAcc-rt
-Llama-8B 79.9 79.9 55.0 54.6 52.1 52.9 3.3 1760.6 54.0 15.7
-Llama-70B 84.6 84.7 72.6 73.1 61.8 62.5 18.7 633.3 52.6 17.9
-GPT-4o-mini 78.4 78.7 55.8 55.8 62.2 62.2 22.8 411.4 51.3 16.8
-GPT-4o 85.0 85.0 76.3 76.5 69.1 69.4 22.9 512.6 59.3 19.5
-O3-mini 85.9 86.1 80.4 80.6 52.1 60.7 25.2 452.1 51.0 17.0
-4.2 Clinical trial design tasks
-Table 3 presents the evaluation of LLMs across five clinical trial design tasks.
-First, arm design appears to be the easiest among the structured design tasks, with all models achieving relatively high accuracy and macro-F1.
-O3-mini achieves the highest scores (85.9 accuracy and 86.1 macro-F1), followed closely by GPT-4o, suggesting that LLMs are generally capable of recovering structured intervention information when explicitly described in the trial protocol.
-In contrast, endpoint design proves to be more challenging.
-All models perform worse on this task, with accuracy ranging from 52.1 to 69.1 and lower macro-F1 across the board.
-The relatively poor performance, even among advanced models like GPT-4o and O3-min.
-Second, sample size estimation remains a particularly difficult task for all models.
-Despite the reasoning-oriented nature of the problem, no model achieves an accuracy above 26%, and mean absolute error (MAE) remains high for all (e.g., 411.4 for GPT-4o-mini and 452.1 for O3-mini).
-This indicates that LLMs, even those designed for reasoning, lack the statistical grounding or calibration mechanisms required to estimate participant counts from trial parameters such as design type, treatment arms, and expected effect size.
-Third, trial completion assessment is also highly challenging.
-Performance on binary outcome prediction (whether a trial is completed or terminated) and termination rationale identification appears near random, with balanced accuracy scores clustered around 50–55% for trial status (BAcc-out) and 15–20% for termination reasoning (BAcc-rt).
-This reflects the difficulty of inferring future feasibility outcomes from trial protocols.
-In a nutshell, these results indicate that current general-purpose LLMs fall short in supporting key trial design decisions.
-While they show some promise in recognizing explicit design elements, they are far less reliable for tasks that require statistical reasoning, feasibility forecasting, or context-aware endpoint selection.
-These gaps highlight the need for further research and specialized model development for clinical trial tasks.
-5 Related work
-Several recent efforts have focused on constructing structured datasets and benchmarks for clinical trials.
-TOP [19] introduced a benchmark for clinical trial outcome prediction with approximately 10K trials, while TrialBench [30] expanded to include 400K trials from ClinicalTrials.gov for predicting trial duration, outcomes, and other attributes.
-CTKG [11] proposed a knowledge graph that captures design entities (e.g., drugs, conditions) from 8K trials, and FinegrainedResults [9] structured 56K trials into arm-level result graphs.
-CTO [14] similarly focused on representing outcome data for 125K interventional trials.
-CARE [10] extracted structured trial results from 700 literature publications, while PubMedKG [31] linked trial identifiers and citations across publications, patents, and trial registries.
-Other lines of work have developed benchmarks for systematic literature review (SLR) tasks.
-TrialReviewBench [28] built a benchmark with 100 systematic reviews for study search, screening, and evidence summarization, and LEADSInstruct [27] scaled these efforts to 21K reviews.
-However, most of these datasets are limited in scope, focusing only on trials’ design or results, lacking ontology alignment, or without accompanying benchmark tasks.
-Our work, TrialPanorama, addresses these limitations by unifying trial records from 16 global sources and linking them to biomedical ontologies.
-It supports both structured design and result representations, and enables a broad suite of benchmark tasks for both trial design and systematic review automation.
-6 Discussion and conclusion
-Despite the broad coverage and structured integration of clinical trial data in TrialPanorama, several limitations should be acknowledged.
-First, the dataset may reflect biases introduced by the data collection process itself.
-For instance, trials registered on ClinicalTrials.gov, primarily based in the United States, are more consistently structured and accessible than those from other countries, potentially leading to overrepresentation of U.S.-centric studies or specific trial designs.
-Similarly, trial information extracted from scientific publications relies on unstructured text, which varies in quality and completeness and may introduce processing bias.
-Second, there is inherent selection bias in the publication landscape, where studies with positive outcomes are more likely to be published, skewing the dataset’s representation of trial attributes and results.
-Third, while LLMs were employed to assist in data extraction and have demonstrated promising accuracy based on sampled validation, systematic evaluation and error mitigation remain areas for future improvement to ensure high data fidelity across the entire dataset.
-Lastly, the benchmark data constructed in this work represents a simplified and task-focused abstraction of real-world clinical trial workflows.
-It does not capture the full complexity and variability of trial design and execution.
-In addition, the selected LLMs and baseline methods evaluated here reflect only a subset of possible approaches; further exploration of model variants, the role of memorization, and prospective evaluation settings will be important for understanding generalizability and real-world utility.
-To summarize, we presented TrialPanorama, a large-scale structured database of clinical trials, along with a benchmark of eight tasks for systematic review and trial design.
-Evaluation of five leading LLMs reveals limited performance, highlighting the need for domain-adapted AI systems in clinical research.
-TrialPanorama provides a foundation to support future development of intelligent tools for clinical trial planning, evidence synthesis, and automation.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: [N/A]
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: [N/A]
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This paper does not involve theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: [N/A]
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: [N/A]
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: [N/A]
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: [N/A]
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: [N/A]
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: [N/A]
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: [N/A]
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: [N/A]
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: [N/A]
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: [N/A]
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: [N/A]
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: [N/A]
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: [N/A]
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
-[…]
-1Contents of Appendix
-Appendix A Database building
-A.1 Raw data collection
+Developing artificial intelligence (AI) for clinical research requires a comprehensive data foundation that supports model training and rigorous evaluation.
+Here, we introduce TrialPanorama, a large-scale structured resource that aggregates 1.6M clinical trial records from fifteen global registries and links them with biomedical ontologies and associated literature.
+To demonstrate its utility, we build a pipeline that constructs 152K training and testing samples for eight key clinical research tasks.
+Three tasks support systematic review workflows, including study search, study screening, and evidence summarization.
+Five tasks focus on trial design and optimization, including arm design, eligibility criteria design, endpoint selection, sample size estimation, and trial completion assessment and rationalization.
+Benchmarking cutting-edge large language models (LLMs) reveals that generic LLMs have limited capability in clinical reasoning.
+In contrast, an 8B LLM we developed on TrialPanorama using supervised finetuning and reinforcement learning wins over the 70B generic counterparts in all eight tasks, with a relative improvement of 73.7%, 67.6%, 38.4%, 37.8%, 26.5%, 20.7%, 20.0%, 18.1%, and 5.2%, respectively.
+We envision TrialPanorama provides a solid foundation for future scaling of AI for clinical research.
+Introduction
+Developing new drugs remains a lengthy and expensive process that often spans close to a decade and costs more than two billion dollars [martin2017much].
+A major share of this burden arises during the human study phase when clinical trials are conducted to evaluate safety and efficacy across phases I through IV.
+Beyond the execution of trials, clinicians, researchers, and trial designers routinely search for and review clinical trial protocols and results.
+These activities support essential objectives such as planning new studies, assessing treatments, and informing clinical decision making in evidence-based medicine [field2010meta, concato2017randomized].
+There is growing interest in using artificial intelligence (AI), and especially large language models (LLMs), to improve the development and evaluation of clinical trials.
+Recent advances have explored a diverse set of tasks including predicting clinical trial outcomes [jin2020predicting, fu2022hint, wang2023spot], matching patients to appropriate trials [jin2024matching, wornow2025zero], extracting real world cohort data from eligibility criteria [yuan2019criteria2query, park2024criteria2query], drafting clinical trial documents [wang2023autotrial, wang2025informgen], and accelerating systematic review of clinical studies [wang2025foundation, wang2024accelerating].
+Nonetheless, most depend on specialized data sources assembled from various specialized repositories of clinical trials.
+As a result, progress is constrained by the absence of standardized and comprehensive clinical trial resources suitable for AI development.
+In this work, we introduce TrialPanorama, a clinical trial database and benchmark suite created from over one million records and fifteen global sources, to provide a broad foundation for trial design, trial review, and the development of AI for clinical trial tasks.
+We assemble clinical trial records from major global registries, including ClinicalTrials.gov [clinicaltrialsgov], the International Clinical Trials Registry Platform [whoictrp], the European Clinical Trials Register [euclinicaltrial], and several national registries [anzctr, chinesechictr].
+We further incorporate trial-related evidence from literature sources such as PubMed.
+Prior studies have curated selected portions of these resources for focused objectives such as extracting trial results [shi2024constructing, naik2024care], annotating drugs and conditions [chen2022knowledge, brbic2024predicting], linking publications to registered trials [xu2020building], and generating structured outcome data [gao2024automatically].
+Here, we put them into a unified representation that integrates these diverse elements, which can substantially strengthen evidence synthesis in systematic reviews and support data driven approaches to future trial planning [dunn2020time, holzhauer2023super, kim2017broadening].
+Specifically, TrialPanorama is created as both a database and a tool suite for extracting clinical trial information and producing benchmark data for AI development in clinical trial applications.
+To illustrate its utility, we present a pipeline that extracts training and testing data from the resource and supports eight trial-related tasks.
+Five tasks focus on trial design and optimization, including arm design, eligibility criteria design, endpoint design, sample size estimation, and trial completion reasoning and prediction.
+Three tasks support systematic literature review workflows, including study search, study screening, and evidence question answering.
+We further evaluate open source and proprietary LLMs on these tasks to establish baseline performance and to demonstrate the value of TrialPanorama for future research.
+Our benchmark further demonstrates the limitations of generic LLMs when applied to clinical trial tasks.
+We find that widely used LLMs display substantial gaps in study search, study screening, evidence summarization, and several trial design and optimization tasks.
+To explore the value of domain-specific data, we fine-tune two open-source models, Llama-3  [dubey2024llama] and Qwen-3  [yang2025qwen3], using training data extracted from TrialPanorama.
+These tuned models achieve competitive or superior performance relative to larger generic LLMs across both systematic review and trial design tasks.
+This improvement highlights the importance of high-quality trial-specific data foundations and demonstrates that TrialPanorama can meaningfully strengthen the capabilities of LLMs developed for clinical trial applications.
+Results
+TrialPanorama: a unified database and knowledge graph of clinical trials
+Public information on clinical trials is available mainly through research publications and registry sites.
+For publications, we collected records from PubMed by focusing on papers labeled as randomized controlled trials and related tags, then applied a machine learning classifier to identify those that truly report clinical trial studies.
+For registry sources, we gathered data from major countries and regions worldwide, including the large registry clinicaltrials.gov and others such as CHICTR in China.
+Together, these sources yield more than 1.6 million trial records.
+As shown in Figure 1b, our final corpus includes 48.7% from PubMed, 31.6% from clinicaltrials.gov, and 19.6% from additional global registries.
+In addition, we collect the systematic literature review publications from PubMed and parse their references to obtain the PubMed ID of the cited clinical studies.
+Based on the sections of the references, we can further get the included and excluded studies from the review.
+As shown in Figure 1a, for PubMed-sourced records, we gather abstracts and full texts when available, then use a large language model to extract core study attributes, including drugs, conditions, phase, primary endpoints, and related information.
+For global registry sources, we implement structured parsers to extract comparable fields so that all records can be transformed into a unified set of study attributes.
+Using this unified data, we apply a sequence of post-processing steps to standardize key fields.
+These include drugs, conditions, adverse events, biomarkers, and outcome measurements, which are normalized through entity extraction and matching with established databases such as DrugBank [drugbank], MedDRA [meddra], TheMarker [10.1093/nar/gkad862], RxNorm [RxNorm], MeSH term [meshterm], and COMET [kirkham2022core].
+The entire pipeline integrates heuristic rules with LLM-based extraction.
+As illustrated in Figure 1a, the goal of the post-processing stage is to construct a unified database of clinical trials.
+This includes a knowledge graph that links trials through shared target drugs, conditions, and other standardized entities, as well as connections based on shared systematic reviews, which often signal domain-relatedness.
+We also create a standardized relational database with multiple tables that capture detailed trial information, including studies, conditions, biomarkers, endpoints, adverse events, disposition, outcomes, and results.
+As shown in Table 1, we compare TrialPanorama against the most related databases from the literature and highlight its comprehensiveness in sources and the data field coverage.
+Figures 1c and 1d show the distributions of target drugs and conditions in TrialPanorama.
+We observe that a small set of agents appears with very high frequency, such as ethanol, ethinyloestradiol, paclitaxel, acetylsalicylic acid, and metformin, while the long tail consists of many agents used in only a modest number of trials.
+A similar pattern is present for conditions, where general categories related to pathologic processes, signs and symptoms, and broad disease classes appear most often, followed by a long tail of more specific disease groups.
+This unified resource enables efficient large-scale extraction and analysis of clinical trial information and serves as a strong data foundation for developing large language models that support clinical research tasks.
+A pipeline for data extraction and LLM development
+TrialPanorama enables efficient extraction of development data for large language models.
+Figure 2a illustrates an example workflow for constructing development data for the sample size estimation task.
+In Step 1, we identify the study fields that serve as inputs to the task.
+For this example, we use study-level attributes such as study type, year, phase, condition, and related descriptors as the trial setup, and use the field target accrual as the sample size value to be predicted.
+We further select the studies with full text so as to extract the statistical assumptions stated in the statistical analysis plan section.
+In Step 2, we organize these inputs into a standardized format for model development, which includes the instruction that specifies the task, the input that provides the trial setup as context, and the answer that contains the ground truth value for supervised learning.
+Through this approach, we can flexibly select study fields from the relational database and the knowledge graph to create instruction data tailored to many clinical research tasks.
+As shown in Figures 2b and 2d, we use the TrialPanorama pipeline to extract development data for three systematic review tasks and five clinical trial design tasks.
+The systematic review tasks cover the core steps of a review workflow.
+The study search task takes a review topic as input and identifies the most relevant clinical trials from literature sources such as PubMed.
+The study screening task takes the review selection criteria and a candidate study set and evaluates the eligibility of each study for inclusion.
+The evidence summarization task takes the results of multiple clinical trials and produces a concise evidence summary that can support future clinical decision-making.
+For clinical trial design tasks, we include four tasks focused on study design fields: sample size estimation, criteria design, endpoint design, and arm design.
+We also include a trial completion assessment task that takes the trial setup as input, predicts whether the trial is likely to complete or terminate, and explains the factors that drive the prediction.
+Together, these tasks establish a comprehensive benchmark for large language models (LLMs) in clinical research settings and provide train-test splits for building vertical models tailored to clinical research applications.
+We hold out the latest studies as the test set to avoid the data leakage issues.
+In Figure 2a, Step 3 applies a combination of supervised fine-tuning (SFT) and reinforcement learning with verifiable reward (RLVR) to adapt generic large language models to clinical research tasks.
+We first train models with SFT for all tasks, then apply RLVR to tasks that require deeper reasoning and offer a verifiable reward signal, such as sample size estimation.
+This staged approach enables broad task coverage while strengthening reasoning ability.
+Building on this framework, TrialPanorama benchmark provides a unified testbed to evaluate both open and proprietary models and to develop domain-adapted variants.
+For example, we adapt Qwen3-8B to clinical research by training it on the TrialPanorama corpus, producing Qwen3-8B-TP.
+As shown in Figure 2c, Qwen3-8B-TP delivers clear gains over the base Qwen3-8B across all tasks.
+Remarkably, this 8B vertical model surpasses Llama-70B, despite using about ten times fewer parameters, and also outperforms GPT-4o, underscoring the value of targeted post-training for clinical research workflows.
+LLMs for clinical trial design and optimization
+We present all results for trial design tasks in Figure 3.
+To benchmark LLM performance on these tasks, we include open source models Qwen3-8B [yang2025qwen3], Llama3-8B, and Llama3-70B [dubey2024llama], together with proprietary models GPT-4o-mini, GPT-4o, and the reasoning model o3-mini.
+For comparison, we also develop domain-adapted LLMs trained on TrialPanorama using the Qwen3-8B and Llama3-8B bases, which yield Qwen3-8B-TP and Llama3-8B-TP, where “TP” is short for TrialPanorama.
+Figures 3a and 3b illustrate the construction of the three trial design tasks.
+Each task presents the target trial setup on the left, including the disease area, treatment modalities, and study objective, followed by a multiple-choice question that mirrors common decisions made during protocol development.
+For arm design, models must determine which treatment arms best match the target trial, considering both the number of arms and the treatment composition.
+For eligibility criteria design, models must choose the inclusion and exclusion criteria that are most appropriate for the stated objective and patient population.
+For endpoint design, models must select the endpoint option that aligns with the clinical goal of the trial.
+Figures 3c show the results of the three trial design tasks.
+Across arm design, eligibility criteria design, and endpoint design, proprietary frontier models outperform the open source baselines, while the domain-adapted models achieve the strongest overall accuracy.
+For arm design, the best open source model reaches 85.0%, whereas GPT-4o and O3-mini reach 86.3% and 89.0%, respectively.
+The domain-adapted models further improve performance, achieving 89.0% for Qwen3-8B-TP and 90.0% for Llama3-8B-TP.
+A similar pattern is observed in eligibility criteria design, where the baseline models plateau between 55.0% and 72.6%, while O3-mini reaches 84.4%.
+Domain adaptation again yields the highest accuracy, with Qwen3 8B TP at 87.1%.
+Endpoint design shows the largest gap between open source baselines and specialized models.
+Baseline models remain below 62.0%, whereas O3-mini reaches 69.1%.
+The domain-adapted variants achieve 74.3% and 78.2%, respectively, demonstrating consistent gains from tailoring the models to the clinical trial domain.
+These results indicate that domain-adapted LLMs provide the most reliable performance across diverse components of trial design.
+Figure 3d shows the construction of the sample size estimation task, which requires multi-step statistical reasoning rather than pattern matching.
+The model must interpret a structured trial synopsis that includes the group definitions for control and experimental arms, along with the statistical assumptions needed for power analysis.
+Given this setup, the model must compute the required sample size that satisfies the primary power requirement.
+We assess model performance by computing the mean absolute error (MAE) of the model-estimated sample size against the original trial’s target accrual, which was made by human statisticians.
+As shown in Figure 3e, open source models exhibit large estimation errors, with mean absolute error ranging from 575.4 to 1769.6.
+Proprietary frontier models reduce the error substantially, reaching 411.4 for GPT-4o-mini and 452.1 for GPT-4o.
+Domain-adapted models achieve the best accuracy overall, with a mean absolute error of 356.3 for Llama3-8B-TP and 393.6 for Qwen3-8B-TP.
+These results highlight both the difficulty of this reasoning-heavy task and the effectiveness of targeted domain adaptation for improving statistical modeling performance.
+Figure 3f shows the trial completion assessment task, where the model predicts whether a clinical trial will terminate early and, if so, selects the most plausible reason among options such as enrollment issues, safety concerns, lack of efficacy, or operational problems.
+The input includes structured trial attributes such as study type, phase, allocation, intervention model, eligibility criteria, outcome measurements, and arm design, requiring the model to integrate heterogeneous signals before making a binary prediction and generating a rationalization.
+As shown in the left panel, open source models achieve balanced accuracy between 52.2% and 56.4%, while GPT-4o reaches 59.3%.
+Domain-adapted models again deliver the highest performance, with Llama3-8B-TP reaching 74.3% and Qwen3-8B-TP reaching 72.8%.
+The right panel reports the rationalization accuracy, which is substantially more challenging.
+Baseline models remain below 26.0%, whereas the domain-adapted variants reach 26.0% for Llama3-8B-TP and 31.1% for Qwen3-8B-TP.
+These results indicate that domain adaptation improves both the termination prediction and the ability to identify the underlying reason, even though rationalization remains a challenging task.
+LLMs for systematic literature review of clinical trials
+Systematic literature review (SLR) is a foundational activity in clinical research, yet it remains one of the most labor-intensive and time-consuming stages of the trial lifecycle.
+In early trial planning, SLR supports evidence-grounded decisions by mapping the landscape of prior studies, identifying safety signals and adverse events in relevant patient groups, and establishing realistic benchmarks for study design.
+In later stages, SLR enables the integration of clinical evidence across historical trials to inform guideline development and clinical practice.
+To capture these essential workflows, we design a set of SLR tasks centered on three key steps: study search, study screening, and evidence summarization.
+These tasks, illustrated in Figure 4, provide a structured benchmark for evaluating how well LLMs can assist researchers throughout the clinical evidence synthesis pipeline.
+Figure 4a illustrates the setup of the study search task.
+Given a review topic specified through the PICO framework, including the target population, intervention, comparison group, and outcome measures, the model must generate an effective search query for a literature database such as PubMed in order to retrieve as many relevant clinical studies as possible.
+We evaluate retrieval performance using Recall@100 and Recall@500, which measure the coverage of ground truth studies (i.e., the ones that were included in the review paper we reference in the test case) within the top retrieved results.
+As shown in Figure 4b, proprietary models and domain-adapted models achieve markedly stronger coverage than general open source baselines.
+GPT-4o and GPT-4o mini outperform Qwen3 and Llama models at both retrieval depths, and the specialized models trained on our dataset Qwen3-8B-TP and Llama3-8B-TP provide further gains.
+Under Recall@100, Qwen3-8B-TP reaches approximately 32% and Llama3-8B-TP reaches approximately 31%, compared with approximately 17% for Qwen3-8B and approximately 19% for Llama3-8B.
+Under Recall@500, Qwen3-8B-TP attains approximately 57%, and Llama3-8B-TP attains approximately 54%, while the strongest proprietary baseline GPT-4o reaches approximately 47%.
+These results show that targeted post-training on our clinical trial corpus substantially enhances an LLM’s ability to generate precise and comprehensive search queries for SLR workflows.
+Figure 4c presents the setup of the study screening task.
+Given a set of papers retrieved for a review topic and the associated inclusion and exclusion criteria specified by the review, the model must decide for each paper whether it should be included or excluded.
+This task reflects the core filtering stage of an SLR workflow, where researchers determine which studies meet the methodological and clinical requirements of the review.
+We evaluate model performance using precision and recall, which together capture how accurately a model identifies eligible studies without introducing excessive false positives or false negatives.
+As shown in Figure 4d, proprietary models again surpass general-purpose open source LLMs, and the domain-adapted models yield the strongest overall performance.
+O3-mini and GPT-4o show clear gains over Qwen3 and Llama baselines, but the TP variants provide the largest improvements.
+Llama3-8B-TP achieves a precision of approximately 72%, and Qwen3-8B-TP reaches approximately 69%, compared with approximately 55% to 60% for the base open source models.
+In terms of recall, Qwen3-8B-TP attains approximately 86%, and Llama3-8B-TP reaches approximately 72%, again exceeding both open source baselines and proprietary models.
+These results indicate that targeted post-training substantially improves an LLM’s ability to apply complex inclusion and exclusion rules and reliably identify relevant studies for SLR workflows.
+Figure 4e shows the setup of the evidence summarization task.
+Given a review topic specified through the PICO framework and the set of studies included after screening, the model is presented with multiple candidate evidence statements and must identify which one accurately reflects the clinical finding, or determine that none of the candidates is correct.
+This formulation mirrors the synthesis stage in SLR workflows, where reviewers must discern which textual statements faithfully represent the aggregated evidence.
+We evaluate model performance using accuracy and macro F1, which jointly capture the ability to select the correct evidence statement while avoiding false selections when no option is appropriate.
+As shown in Figure 4f, proprietary models outperform general open source baselines, and domain-adapted models again achieve the strongest results.
+GPT-4o and O3-mini provide clear improvements over Qwen3 and Llama models, but the TP variants show the most substantial gains.
+Qwen3-8B-TP reaches an accuracy of approximately 82% and a macro F1 of approximately 81%, the highest among all evaluated models.
+Llama3-8B-TP also performs well with accuracy near 76% and macro F1 near 75%, compared with accuracy values between 67% and 74% for the general-purpose baselines.
+These results demonstrate that targeted post-training on clinical trial data significantly enhances an LLM’s ability to distinguish accurate evidence from distractors, a critical capability for reliable evidence synthesis in SLR workflows.
+Discussion
+TrialPanorama represents a step toward enabling cutting-edge large language models (LLMs) to support clinical research, an essential component in accelerating therapy development and improving population health.
+Our work addresses two central challenges in building effective domain models for this setting.
+First, developing domain-specialized LLMs requires high-quality and large-volume vertical data, yet such resources have historically been fragmented or unavailable at scale.
+Second, there has been no standardized pipeline for transforming raw clinical research data into development-ready training corpora or for applying post-training strategies, including supervised finetuning (SFT) and reinforcement learning with verifiable reward (RLVR), that reliably strengthen model reasoning.
+Here, we first assemble over 1.6 million clinical trial records and 30K systematic literature reviews from literature databases, ClinicalTrials.gov, and additional global trial registries.
+After multi-stage cleaning, harmonization, and ontology-based standardization, these sources are integrated into TrialPanorama, which provides a unified trial database and a knowledge graph linking review papers to clinical studies and connecting studies through shared drugs, conditions, and other biomedical entities.
+Using this resource, we show how developers can efficiently extract structured supervision and construct development data for training and evaluating LLMs across five trial design and optimization tasks and three systematic review tasks.
+Our benchmark compares eight open and proprietary models alongside TrialPanorama adapted LLMs.
+The results consistently demonstrate that domain-adapted models outperform all generic baselines.
+Through the combination of TrialPanorama and a unified SFT plus RLVR pipeline, we observe substantial gains in reasoning quality.
+In particular, our 8B-level models surpass a 70B generic model and a proprietary reasoning model such as o3-mini, underscoring that task-aligned data and verifiable reward signals can be more consequential than scale alone for clinical research reasoning.
+This study has several limitations.
+First, although TrialPanorama integrates one million trials from fifteen global registries, coverage remains uneven, particularly for regions with limited reporting, and the resulting representativity may vary across therapeutic areas and countries.
+The pipeline also depends heavily on registry-level text quality for fields such as eligibility criteria and outcome descriptions.
+Since large-scale processing relies on a mixture of heuristics and AI-based extraction, low-quality or ambiguous text can introduce noise into task inputs or training targets.
+Ontology normalization mitigates variability but cannot fully resolve missing or underspecified content.
+Second, the evaluation setup necessarily simplifies real-world clinical research workflows.
+We adopt proxy metrics and task formulations to enable reproducible benchmarking, but these abstractions cannot capture the full complexity of clinical decision making or the nuances of protocol design.
+Similarly, post-training approaches such as RLVR are applicable only to tasks with clearly verifiable labels, leaving many clinically important tasks outside the current scope.
+Third, domain generalizability and temporal drift remain challenges.
+Registry data reflect design practices at the time of registration, which may not align with evolving standards of care or contemporary regulatory expectations.
+Without continual updates, models trained on historical patterns may not fully capture emerging paradigms such as decentralized, adaptive, or platform trials.
+Although TrialPanorama covers a broad range of conditions and study types, model performance may vary in underrepresented therapeutic domains or rare disease contexts.
+In conclusion, this work demonstrates an end-to-end pipeline for developing LLMs tailored to clinical research, spanning data construction, knowledge integration, and model adaptation.
+By building on a large-scale, queryable database and combining supervised finetuning with reinforcement learning using verifiable rewards, we show that domain-aligned data and post-training strategies can meaningfully elevate model performance on core clinical research tasks.
+We envision that such data foundations, together with continued advances in algorithms, will play a central role in enabling reliable and impactful AI systems for biomedicine.
+Figure 1: Overview of TrialPanorama and its end-to-end processing workflow.
+a, Raw clinical research data are collected from fifteen global sources, including PubMed, ClinicalTrials.gov, and international registries.
+LLM-based extraction, hybrid rule-based methods, and ontology-grounded normalization are applied to derive unified study attributes, followed by entity extraction and matching to DrugBank, MedDRA, RxNorm, MeSH, and other biomedical ontologies.
+The processed information is stored in both a relational schema and a heterogeneous knowledge graph linking trials, publications, drugs, diseases, and outcomes.
+b, Distribution of data sources in TrialPanorama showing contributions from PubMed, ClinicalTrials.gov, and aggregated global registries, together with a country-level breakdown of non-US registries.
+c, Most frequent drug entities mapped to standardized identifiers, illustrating the long tail distribution and coverage of therapeutic compounds.
+d, Most frequent disease and condition entities after ontology harmonization, demonstrating broad coverage across oncology, infectious disease, neurologic disease, metabolic disorders, cardiovascular disease, and other major therapeutic areas.
+Figure 2: Benchmark construction and model performance across systematic review and clinical trial design tasks.
+a, End-to-end workflow for transforming structured and semi-structured clinical trial records into supervised training examples.
+Extracted study attributes are composed into standardized instruction input answer triples, enabling the development of models for tasks such as sample size estimation, criteria design, and endpoint design.
+b, Overview of the benchmark for systematic review tasks, including study search, study screening, and evidence summarization, together with the scale of training and testing sets derived from real-world review corpora.
+c, Overall performance of representative large language models on the full suite of benchmark tasks.
+“Qwen3-8B-TP*” indicates Qwen3-8B model fine-tuned using the mixture of supervised fine-tuning (SFT) and reinforcement learning with verifiable reward (RLVR) strategies on TrialPanorama.
+Results illustrate substantial variation across tasks, with domain-adapted models achieving stronger accuracy and recall in both review-oriented and design-oriented evaluations.
+d, Benchmark tasks for clinical trial design covering sample size estimation, eligibility criteria design, endpoint selection, arm design, and trial completion assessment, along with the number of training and testing instances for each task.
+Figure 3: Benchmark tasks for clinical trial design and comparative performance of large language models.
+a, Example trial setup illustrating how key attributes, such as indication, interventions, and study objective, form the context for downstream design tasks.
+b, Illustrations of three multiple-choice design tasks: arm design, eligibility criteria design, and endpoint design, where the model must select the option that best aligns with the target trial.
+c, Accuracy of eight models on arm design, eligibility criteria design, and endpoint design, showing clear gains from domain-adapted training.
+d, Illustration of the sample size estimation task, including group definitions and statistical assumptions used to construct ground truth labels.
+e, Mean absolute error for sample size estimation across eight models, where lower error indicates better performance.
+f, Trial completion assessment task and corresponding performance, evaluated by balanced accuracy for termination prediction and rationale selection, highlighting improvements in models trained on TrialPanorama.
+Figure 4: Systematic review task setup and model performance across study search, study screening, and evidence summarization.
+a, Study search task formulation: given a review topic specified by the PICO elements, the model must generate an effective PubMed query to retrieve relevant clinical studies.
+b, Performance of eight models on study search measured by Recall at one hundred and Recall at five hundred, highlighting substantial gains from domain-adapted training.
+c, Study screening task formulation: given review selection criteria and candidate studies, the model classifies each study as included or excluded.
+d, Precision and recall for study screening across models, showing improvements in both specificity and sensitivity with domain-adapted models.
+e, Evidence summarization task formulation: given a review topic and selected studies, the model must identify the correct evidence statement among multiple candidates or select none if no option is supported.
+f, Accuracy and Macro F1 for evidence summarization, demonstrating strong benefits of task-aligned training for fine-grained evidence reasoning.
+Table 1: Comparison of TrialPanorama with prior clinical trial datasets and benchmarks.
+The table summarizes sample size, data sources, and coverage of key structured components, including design attributes, trial results, clinical outcomes, study publication relations, and ontology linkage.
+It also highlights whether each resource provides benchmark tasks for trial design or systematic review workflows, showing that TrialPanorama offers the most comprehensive coverage across all dimensions.
+Datasets Sample size Data source Design Result Outcome Relation Ontology Design Benchmark Review Benchmark
+TrialPanorama (Ours) 1.6M Multiple (15 sources)
+TOP [fu2022hint] 10K ClinicalTrials.gov
+CTKG [chen2022knowledge] 8K ClinicalTrials.gov
+ClinicalTrials.govResults [shi2024constructing] 56K ClinicalTrials.gov
+CTO [gao2024automatically] 125K ClinicalTrials.gov
+TrialBench [chen2024trialbench] 400K ClinicalTrials.gov
+CARE [naik2024care] 700 pubs PubMed
+PubMedKG [xu2024pubmed] 480K PubMed,ClinicalTrials.gov
+TrialReviewBench [wang2024accelerating] 100 reviews PubMed,ClinicalTrials.gov
+LEADSInstruct [wang2025foundation] 21K reviews PubMed,ClinicalTrials.gov
+Method
+Data collection
 To construct our benchmark dataset, we crawled clinical trial and publication records from a diverse set of publicly accessible registries and databases.
-The data URLs and the date of the most recent crawl are listed in Table 4.
-Table 4: List of data sources and the corresponding last crawl dates.
-Source Data URL Last Crawled
-ANZCTR https://www.anzctr.org.au/ April 1, 2025
-Brazilian Registry https://ensaiosclinicos.gov.br April 1, 2025
-ChiCTR https://www.chictr.org.cn/searchprojEN.html April 1, 2025
-Dutch Registry https://www.onderzoekmetmensen.nl/en/ April 1, 2025
-EU Clinical Trials https://www.clinicaltrialsregister.eu April 1, 2025
-DRKS https://drks.de/search/en/results April 1, 2025
-Iranian Registry https://irct.behdasht.gov.ir/ April 1, 2025
-ISRCTN https://www.isrctn.com April 1, 2025
-Japan Registry https://center6.umin.ac.jp/cgi-open-bin/ctr_e/ctr_view.cgi April 1, 2025
-South Korea CRIS https://cris.nih.go.kr/cris/index/index.do April 8, 2025
-Pan African Registry https://pactr.samrc.ac.za/Search_v2.aspx April 8, 2025
-Sri Lanka CTR https://slctr.lk/trials April 8, 2025
-Thai Registry https://www.thaiclinicaltrials.org April 8, 2025
-REec https://reec.aemps.es/reec/public/web.html April 9, 2025
-CTRI https://ctri.nic.in/Clinicaltrials/advancesearchmain.php April 9, 2025
-PubMed https://ftp.ncbi.nlm.nih.gov/ April 9, 2025
-After collection, the raw data was subjected to systematic cleaning procedures to ensure consistency and reliability across sources:
-  • Intra-source de-duplication: Duplicate entries within the same registry were detected and removed using unique identifiers (such as trial registration numbers), and corroborated using additional metadata like titles.
-  • Inter-source de-duplication: Clinical trials registered in multiple registries were identified via fuzzy matching on titles. Duplicate records across sources were merged into a single one.
-  • Field normalization: Metadata fields such as trial status, registration date, and condition labels were standardized across sources to ensure consistent querying.
-  • Missing value handling: Entries with missing or incomplete critical fields were flagged.
+The data URLs and the date of the most recent crawl are listed in Supplementary Table 1.
+After collection, the raw data were subjected to systematic cleaning procedures to ensure consistency and reliability across sources: (1) Intra-source de-duplication: Duplicate entries within the same registry were detected and removed using unique identifiers (such as trial registration numbers), and corroborated using additional metadata like titles.
+(2) Inter-source de-duplication: Clinical trials registered in multiple registries were identified via fuzzy matching on titles.
+Duplicate records across sources were merged into a single one.
+(3) Field normalization: Metadata fields such as trial status, registration date, and condition labels were standardized across sources to ensure consistent querying.
+(4) Missing value handling: Entries with missing or incomplete critical fields were flagged.
 These operations minimized noise, reduced redundancy, and created a unified and clean dataset suitable for benchmarking tasks.
-A.2 Studies table
+Database building
+TrialPanorama is organized into a set of structured tables that together provide a unified view of global clinical research data.
+The Studies table serves as the central entry for each trial and aggregates identifiers from registries and publications, trial design attributes such as phase and sponsor, and primary text fields including titles and summaries.
+Several complementary tables enrich each study with standardized clinical entities.
+The Conditions table stores disease terms mapped to Medical Subject Headings, including both primary concepts and hierarchical ancestors.
+The Drugs table provides a normalized representation of pharmacologic interventions, including cleaned drug names, RxNorm and DrugBank identifiers, and approval status from major regulatory agencies.
+The Disposition table captures the arm structure of each trial by pairing treatment groups with intervention descriptions, while the Adverse Events table records arm-specific safety findings from ClinicalTrials.gov, including frequencies, seriousness categories, and MedDRA codes.
+To support evidence extraction and comparative analysis, we constructed additional tables centered on trial findings and clinical endpoints.
+The Trial Results table provides PICO structured summaries of outcome data at the arm level, combining extractions from PubMed abstracts with structured results from ClinicalTrials.gov.
+The Trial Outcomes table assigns unified outcome types to completed and terminated studies, including positive and negative classifications and fine-grained termination reasons.
+The Endpoints table standardizes primary outcomes by mapping them to COMET domains and subdomains, creating a consistent representation of clinical endpoints across heterogeneous sources.
+The Biomarkers table links text spans from trial descriptions to curated therapeutic biomarkers from TheMarker and records biomarker name, functional class, biomarker type, and associated genes when applicable.
+Finally, the Relations table encodes connections among systematic reviews, cited studies, and registry records.
+Each entry is a structured triple that captures include, exclude, awaiting assessment, or cross registry link relationships.
+Together, these tables form a comprehensive relational database that supports large-scale analysis of trial design, evidence synthesis, and clinical development patterns.
+Detailed procedures for extraction, normalization, and validation are reported in Supplementary Section 1.
+Trial design tasks: arm design
+Definition & Motivation.
+This task assesses whether the model can identify the correct treatment arms and intervention details for a given clinical trial.
+Based on the trial’s title and brief summary, the model must choose among four candidate descriptions the one that accurately represents the trial’s arm labels, arm types, and assigned interventions.
+An example instance is shown in Supplementary Table 8.
+Input & Output.
+The model is provided the official trial title and summary, followed by four arm descriptions that list the group labels, intervention types, and associated treatments.
+It must return the label corresponding to the correct description.
+Dataset Construction.
+We start with sets of up to four trial identifiers drawn from trials cited together in the same Cochrane review.
+For each identifier, a script extracts the arm groups and individual intervention details from our trial metadata.
+If fewer than four candidates are available, additional trials are added at random (with a fixed seed for reproducibility).
+To create challenging distractors, incorrect descriptions are drawn from the same arm and intervention information of other trials cited by that review.
+All items are then sorted by the numeric portion of the trial identifier and partitioned so that the latest 1,000 trials form the test set, the next 500 form validation, and the rest form training.
+Evaluation Metrics.
+We report accuracy and macro-F1 to account for class imbalance across answer choices.
+Trial design tasks: eligibility criteria design
+Definition & Motivation.
+This task measures the model’s ability to match a trial’s inclusion and exclusion criteria.
+Given a trial’s title and summary, the model must select, from four text options, the one that accurately lists the trial’s eligibility rules.
+An example instance is shown in Supplementary Table 9.
+Input & Output.
+The model receives the trial title and summary along with four candidate eligibility criterion sets. It must output the correct option label.
+Dataset Construction.
+For eligibility criteria, dataset construction follows the same approach as in Arm Design: we form candidate sets from trials cited by the same Cochrane review, extract each trial’s eligibility rules from our metadata, pad with random trials when needed, and draw distractors from other cited trials to ensure relevance.
+Examples are then sorted by the numeric part of the trial identifier and split into test (latest 1,000), validation (next 500), and training (remainder).
+Evaluation Metrics.
+We report Accuracy and macro‑F1 to account for class imbalance across answer choices.
+Trial design tasks: endpoint design
+Definition & Motivation.
+This task evaluates whether the model can identify the primary outcome measures of a clinical trial.
+Based on the trial title and summary, the model must pick the option that lists the correct primary endpoints.
+An example instance is shown in Supplementary Table 10.
+Input & Output.
+The input consists of the trial title and summary plus four outcome descriptions that concatenate the primary measures and any associated timeframes. The model must choose the correct label.
+Dataset Construction.
+For primary endpoints, dataset construction also follows the Arm Design template: trials cited together provide candidate sets, we extract each trial’s primary outcomes from our metadata, pad with random examples if needed, and sample distractor options from other cited trials.
+The resulting items are ordered by the numeric portion of the trial identifier and divided into test (latest 1,000), validation (next 500), and training splits.
+Evaluation Metrics.
+We report Accuracy and macro‑F1 to account for class imbalance across answer choices.
+Trial design tasks: sample size estimation
+Definition & Motivation.
+This task evaluates the model’s quantitative reasoning by asking it to estimate the number of participants required for a clinical trial based on its design and stated statistical assumptions.
+Success requires understanding trial objectives, translating textual assumptions (effect size, significance level, power, dropout) into a numerical enrollment target, and applying standard sample‐size formulas.
+An example instance is shown in Supplementary Table 7.
+Input & Output.
+As input, the model receives: a narrative description of the trial design—official title, conditions, arm groups, interventions and primary outcomes, as well as a concise summary of the statistical assumptions extracted from the protocol text (treatment effect, alpha, power, expected dropout).
+The model must output a single integer corresponding to the estimated sample size.
+Dataset Construction & Splits.
+We linked 5,737 PubMed protocols to their NCTIDs and identified 4,315 full-text sections titled “sample size” together with 5,678 CT.gov records containing enrollment counts.
+Matching the reported CT.gov enrollment with the corresponding paper section yielded 2 143 examples where the protocol paragraph explicitly mentions the official enrollment number.
+For each example, we used a deterministic prompt to summarize the paragraph into (1) effect‐size assumptions, (2) statistical parameters (significance level, power, dropout rate), and (3) the true enrollment.
+We then generated a question by combining trial metadata with these summarized assumptions.
+Finally, all examples were sorted by the numeric part of the NCTID; the most recent 1,000 form the test set, the next 500 the validation set, and the remainder the training set.
+Evaluation Metrics.
+We report (1) the mean absolute error between predicted and actual enrollments, and (2) the proportion of trials for which the model’s estimate lies within ±20 % of the true sample size.
+Trial design tasks: trial completion assessment
+Definition & Motivation.
+This task simulates a real-world feasibility analysis during clinical trial planning, where the model must assess whether a trial is likely to reach completion or be terminated early based on its protocol design.
+In addition, if a trial is predicted to terminate, the model is asked to identify the most likely reason for termination from a predefined set of categories (e.g., enrollment issues, safety concerns, lack of efficacy, operational problems).
+This reflects practical needs in the early design stage when sponsors assess trial viability.
+An example instance is shown in Supplementary Table 11.
+Input & Output.
+Each example consists of structured metadata and protocol text for a registered interventional study, including trial phase, number of arms, masking, intervention model, primary purpose, arm descriptions, and detailed eligibility criteria.
+The output is a binary label indicating whether the trial was completed or terminated.
+For trials labeled as terminated, the model must also predict the most likely termination reason from one of several standardized categories.
+Dataset Construction.
+We construct the dataset using completed and terminated interventional trials from ClinicalTrials.gov with known outcomes.
+Structured fields are extracted from the studies, disposition, eligibility_criteria, and arms tables in our trial database.
+Only trials with definitive outcome labels (i.e., Completed or Terminated) and non-missing eligibility criteria are retained.
+Termination reasons are normalized into broad categories using a curated taxonomy (e.g., enrollment issues, lack of efficacy, safety, sponsor decision).
+To avoid label leakage, we exclude any outcomes or termination reasons from the input text fields.
+Evaluation Metrics.
+We evaluate trial completion prediction using accuracy. For terminated trials, termination reason classification is evaluated using macro-F1 to balance across categories with different frequencies.
+Systematic literature review tasks: study search
+Definition & Motivation.
+This task simulates the study search step in conducting systematic reviews, where the model must comprehend a given research question and synthesize a corresponding search query to retrieve relevant studies for inclusion.
+The key motivation behind this task is to streamline a traditionally labor-intensive process that typically requires the expertise of trained librarians, particularly due to the complexity and variability of search query formats across literature databases.
+Automating this step can significantly accelerate systematic review workflows.
+An example instance is presented in Supplementary Table 4.
+Input & Output.
+The input for this task includes the setup of a systematic review, consisting of the research background, objective, and study selection criteria.
+The model is expected to generate a structured search query, which is then submitted to the PubMed Search API to retrieve the top 500 results and their associated PubMed IDs.
+A successful output should go beyond simple keyword extraction, incorporating semantic understanding and augmentation of the input to produce a comprehensive and effective query.
+Dataset Construction.
+We constructed the dataset using a relation table that links systematic reviews to primary studies.
+Specifically, we focused on instances with the relation type include, indicating that the study was selected for inclusion in the review.
+For each review, we parsed its full text to extract relevant sections and used the linked PubMed abstracts as ground-truth targets.
+This process resulted in a dataset of 8,104 examples derived from systematic review articles, covering more than 180,000 included PubMed studies.
+We reserved the most recent 100 reviews as a held-out test set.
+Evaluation Metrics.
+We evaluate this task as a retrieval problem by measuring how many ground-truth studies appear in the top search results produced by the model-generated query.
+Specifically, we report Recall@100 and Recall@500, reflecting the number of included studies retrieved among the top 100 and 500 PubMed results, respectively.
+Systematic literature review tasks: study screening
+Definition & Motivation.
+This task simulates the study selection step in a systematic review pipeline, where the model must assess the eligibility of candidate studies for inclusion based on the review’s predefined criteria.
+Given a systematic review’s background, objective, and eligibility criteria, along with a set of candidate studies retrieved from PubMed (each with a title and structured abstract), the model determines whether each study should be included or excluded.
+This step is often labor-intensive and requires detailed reasoning across clinical nuances and protocol designs.
+An example instance is shown in Supplementary Table 5.
+Input & Output.
+The input to this task consists of the systematic review protocol (comprising the review’s background, objective, and selection criteria) and a list of candidate studies, each represented by its title and structured abstract.
+The output is a binary label (included or excluded) for each candidate study, indicating whether it meets the review’s eligibility criteria.
+Accurate performance requires the model to match eligibility conditions against sometimes subtle or incomplete cues in the candidate study abstracts.
+Dataset Construction.
+We construct the dataset using reviews that have explicit inclusion and exclusion annotations for candidate studies, available from the relations table in our clinical trial knowledge base.
+For each review, we extract its background, objective, and selection criteria from the abstract, and retrieve associated candidate studies along with their labels (included or excluded) based on the relation types include and exclude.
+To ensure a balanced and informative evaluation, we downsample the candidate pool for each review to include a random mix of included and excluded studies, preserving diversity while controlling for class imbalance.
+Each instance contains up to 10 candidate studies, each with a title and structured abstract, paired with a single review protocol.
+Evaluation Metrics.
+We evaluate model performance using three metrics: accuracy, precision, and recall.
+Accuracy reflects the overall proportion of correctly classified studies.
+Precision measures the proportion of correctly identified included studies among all studies predicted as included, while recall measures the proportion of truly included studies that were correctly identified by the model.
+These metrics jointly assess both correctness and completeness of model screening decisions.
+Systematic literature review tasks: evidence summarization
+Definition & Motivation.
+This task simulates the final synthesis step of a systematic review, in which the model must draw a conclusion from the included studies’ findings.
+Given a self-contained multiple-choice question based on a Cochrane review’s objectives, the model selects the answer option that aligns with the review’s main conclusion.
+An example instance is shown in Supplementary Table 6.
+Input & Output.
+The input to the evidence summarization task consists of a self‐contained multiple‐choice question generated from a Cochrane review abstract (including its answer options A, B, C, …) together with the titles and full structured abstracts of all studies included in that review (ranging from 1 to 20 PMIDs).
+The model’s output is a single label (A, B, C, …) corresponding to the correct conclusion.
+Successfully performing this task requires the model to comprehend scientific text, aggregate evidence across multiple documents, and map study findings to the correct conclusion.
+Dataset Construction.
+First, Cochrane review abstracts are processed by GPT-4o using a prompt that instructs the model to generate a self‐contained multiple‐choice question based on the review objectives, to produce answer options covering plausible outcomes, and to output only valid JSON with the keys question, options, and answer.
+The prompt also forbids referencing the original review by phrases such as “based on” to ensure questions are fully self‐contained, and the model is run at temperature 0 to guarantee deterministic outputs.
+Next, for each generated question, we retrieve the included‐study PMIDs from the review metadata and load the corresponding titles and structured abstracts from our PubMed cache, discarding any entries with zero or more than twenty included studies.
+Finally, the remaining $N$ examples are sorted by numeric PMID (serving as a proxy for publication date); the most recent 1,000 examples form the test set, the preceding 500 form the validation set, and the remainder form the training set.
+Evaluation Metrics.
+We report accuracy and macro-F1 to account for class imbalance across answer choices.
+Model development
+We developed two domain-adapted large language models (LLMs) based on LLaMA-3-8B-Instruct [dubey2024llama] and Qwen3-8B [yang2025qwen3], respectively.
+Both models were trained following a two-stage pipeline that combines supervised fine-tuning (SFT) for domain knowledge injection and reinforcement learning with verifiable rewards (RLVR) for improving task-specific performance.
+Supervised Fine-tuning (SFT)
+Supervised fine tuning was performed in two sequential phases.
+In the first phase, the model was trained only on the tasks with high volume datasets, specifically trial design and trial completion assessment.
+In the second phase, the model was trained on the remaining tasks, including sample size estimation, study search, study screening and evidence summarization.
+Separating these phases ensured that high volume tasks did not dominate the gradient signals, otherwise the substantially smaller tasks would struggle to be adequately learned by the model.
+The model was optimized using the next token prediction objective applied only to answer tokens.
+Let $x$ denote the input sequence, which includes both the instruction and the question, and let $y=(y_{1},\dots,y_{K})$ denote the target answer sequence.
+The supervised fine tuning objective is
+$\mathcal{L}_{\mathrm{SFT}}=-\sum_{t=1}^{K}\log p_{\theta}\bigl(y_{t}\mid y_{<t},x\bigr),$ (1)
+so that the input $x$ only serves as conditioning context, while the loss is computed exclusively over the answer tokens $y$.
+We implement SFT recipe by using the LLaMA Factory framework [zheng2024llamafactory].
+Training used a context length of 8192 tokens and a peak learning rate of $1\times 10^{-6}$ under a cosine decay schedule.
+The AdamW optimizer was employed with a batch size of 16 and numerical precision was set to bfloat16.
+To increase computational efficiency, DeepSpeed ZeRO-3 [rajbhandari2020zero] and FlashAttention-2 [dao2024flashattention] were incorporated.
+Phase 1 was trained for three epochs, and Phase 2 continued for an additional ten epochs.
+Reinforcement Learning with Verifiable Rewards (RLVR)
+After SFT, we further optimized the models using RLVR for tasks that admit automatically verifiable outcomes, specifically sample size estimation and study search.
+Training followed a two-step recipe.
+The first step trained the model on sample size estimation, and the second step trained the model on study search.
+RLVR was implemented using the VeRL framework [sheng2025hybridflow] with GRPO [shao2024deepseekmath] as the underlying optimization algorithm.
+For each input $x$, we sample $G$ output sequences $y_{i}=(y_{i,1},\dots,y_{i,|y_{i}|})$ from the old policy $\pi_{\theta_{\mathrm{old}}}(\cdot\mid x)$, and compute a scalar reward $r_{i}$ for each sequence.
+The GRPO objective maximizes a clipped policy improvement term averaged across all rollouts and all tokens:
+$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{x,\;\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid x)}\left[\frac{1}{G|y_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|y_{i}|}\min\!\left(\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(y_{i,t}\mid x,y_{i,<t})}A_{i},\;\operatorname{clip}\!\left(\frac{\pi_{\theta}(y_{i,t}\mid x,y_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(y_{i,t}\mid x,y_{i,<t})},1-\varepsilon,\;1+\varepsilon\right)A_{i}\right)\right]$
+where the scalar advantage $A_{i}$ for rollout $i$ is computed by normalizing rewards within the group of $G$ rollouts,
+$A_{i}=\frac{r_{i}-\mathrm{mean}(\{r_{1},\dots,r_{G}\})}{\mathrm{std}(\{r_{1},\dots,r_{G}\})}.$ (3)
+Training used a prompt length limit of 2,048 tokens, a response length limit of 512 tokens, batch sizes of 16 for training and 32 for validation, and a learning rate of $1\times 10^{-6}$.
+During rollout generation, the sampling temperature was set to 0.6, nucleus sampling used a top-p of 0.95, and eight rollouts were generated per prompt.
+Each RLVR stage was trained for five epochs.
+Across both tasks, rewards combined a format reward and a performance reward.
+The format reward enforced adherence to the required output structure.
+Specifically, the model was required to present its final answer within an <answer> … </answer> tag pair.
+Outputs that failed to satisfy this requirement received a penalty of 2-2.
+When the answer was correctly enclosed within the prescribed tags, the format reward was set to $0.1$.
+Conditional on correct formatting, the performance reward evaluated task accuracy.
+For the study search task, performance was defined by retrieval quality, measured using recall@100@100.
+For the sample size estimation task, performance was determined by whether the predicted sample size fell within a tolerance window between 0.8 and 1.2 times of the ground truth.
+Further, if the performance reward was zero or negative, the formatting bonus was not applied.
+This reward structure ensures that correctly formatted but substantively incorrect outputs do not accumulate positive reward, thereby aligning the learning signal with the practical expectations.
+The final models, LLaMA-3-8B-TP and Qwen-3-8B-TP, were produced by applying the supervised fine-tuning and reinforcement learning pipeline described above.
+These models exhibit substantially improved performance across trial design tasks and systematic review tasks compared with their base versions, indicating the importance of domain aligned post training strategies.
+Data Availability
+The database is available at https://huggingface.co/datasets/zifeng-ai/TrialPanorama-database.
+The built training and evaluation datasets are available at https://huggingface.co/datasets/zifeng-ai/TrialPanorama-benchmark.
+Code Availability
+The code for experiments used in this study is available at https://github.com/RyanWangZf/TrialPanorama.
+Supplementary Notes
+1 Database building
+1.1 Studies table
 There are three main sources of clinical trial records in our dataset: ClinicalTrials.gov, global registries, and PubMed.
@@
 While this method does not capture the full complexity of trial protocols, it provides a scalable way to convert a large volume of unstructured publications into structured records suitable for downstream analysis.
-A.3 Conditions table
+1.2 Conditions table
 To standardize disease conditions associated with each trial, we extracted Medical Subject Headings (MeSH) terms using the NCBO BioPortal Annotator API [^2] https://bioportal.bioontology.org/annotator.
-This process was applied only to non–ClinicalTrials.gov sources (e.g., PubMed, Embase, and international registries), as ClinicalTrials.gov conditions are already standardized.
+This process was applied only to non–ClinicalTrials.gov sources (e.g., PubMed and international registries), as ClinicalTrials.gov conditions are already standardized.
 We used the official title and brief summary of each trial as input to retrieve disease-related MeSH terms (semantic type T047) and their hierarchical ancestors.
 The resulting MeSH types were labeled as mesh-list (primary condition), mesh-ancestor (intermediate ancestors), and mesh-ancestor-main (top-level ancestor).
-A.4 Drugs Table
+1.3 Drugs Table
 To construct a comprehensive drugs table, we extract intervention mentions from three primary sources: (1) ClinicalTrials.gov and (2) PubMed.
@@
 Table 2 summarizes the drug mapping statistics for each data source, including total unique drug mentions, RxNorm mapping counts and rates, and DrugBank resolution results.
-Table 5: Drug mapping statistics across data sources.
-Source Total Drugs RxNorm Mapped RxNorm Rate (%) DrugBank Mapped DrugBank Rate (%)
-PubMed 953,361 579,035 60.74 241,879 25.37
-ClinicalTrials.gov 366,214 290,371 79.29 273,509 74.69
-Supplementary Figure 1: Pipeline for drug name normalization and standardization.
-Drug names are cleaned, mapped to RxNorm, and subsequently resolved to DrugBank identifiers using a combination of indirect and direct strategies.
 To further enrich the standardized drug table, we annotate each drug entry with regulatory approval status.
@@
 Food and Drug Administration (FDA), the European Medicines Agency (EMA), and Japan’s Pharmaceuticals and Medical Devices Agency (PMDA).
-These lists are sourced from DrugCentral 2023 [33] [^3] https://drugcentral.org/, which provides a consolidated view of approval data across jurisdictions.
+These lists are sourced from DrugCentral 2023 [avram2023drugcentral] [^3] https://drugcentral.org/, which provides a consolidated view of approval data across jurisdictions.
 Approval annotations are applied by checking each drug’s normalized name, RxNorm-mapped name, and DrugBank-mapped name against the cleaned lists of approved drug names.
@@
 This process enables us to systematically track the clinical regulatory status of drugs within our dataset, aiding downstream tasks that require distinguishing between approved and experimental interventions.
-A.5 Trial results table
+1.4 Trial results table
 To construct the Trial Result Table, a core component for the evidence summarization benchmark task, we systematically extracted key findings from two primary types of sources within our TrialPanorama database: 1) published clinical trial literature, encompassing abstracts from PubMed, and 2) structured trial data directly from ClinicalTrials.gov registry entries.
@@
 Abstract: {abstract}
-A.6 Trial outcomes table
-Supplementary Figure 2: (a) Trial outcome extraction workflow.
-Completed trials are labeled using outcome measures and matched PubMed abstracts.
-Terminated trials are categorized by mapping stop reasons to predefined outcome types via keyword similarity.
-(b) Distribution of extracted outcome types on trials from ClinicalTrials.gov.
+1.5 Trial outcomes table
 To extract trial outcomes, we developed two distinct workflows for trials registered on ClinicalTrials.gov and those from other international sources.
@@
 We provided the trial title, primary outcome measures, and result text to GPT-4o to infer the trial outcome.
-A.7 Relations table
-Table 6: Statistics of the relation types in the relations table.
-Relation Type Count Percentage (%)
-PubMed - cite - PubMed 349,680 37.177
-PubMed - exclude - PubMed 237,076 25.205
-PubMed - include - PubMed 180,827 19.225
-PubMed - linked_to - ClinicalTrials.gov 167,121 17.768
-ClinicalTrials.gov - linked_to - JAPAN 2,787 0.296
-PubMed - include - ClinicalTrials.gov 954 0.101
-ClinicalTrials.gov - linked_to - KOREA 632 0.067
-PubMed - cite - ClinicalTrials.gov 425 0.045
-PubMed - exclude - ClinicalTrials.gov 406 0.043
-ClinicalTrials.gov - linked_to - GERMAN 331 0.035
-ClinicalTrials.gov - linked_to - ANZCTR 201 0.021
-ClinicalTrials.gov - linked_to - PAN_AFRICAN 56 0.006
-ClinicalTrials.gov - linked_to - SRI_LANKA 37 0.004
-ClinicalTrials.gov - linked_to - CHICTR 31 0.003
-ClinicalTrials.gov - linked_to - DUTCH 18 0.002
+1.6 Relations table
 We implemented a multi-stage data processing pipeline centered on Cochrane Library reviews and their linked PubMed metadata to construct structured relationships between systematic reviews, clinical trials, and associated publications.
@@
 The statistics of all the relation types are in Supplementary Table 3.
-A.8 Dispositions table
+1.7 Disposition table
 Disposition reflects an essential aspect of a clinical trial: the design of experimental arms, including details such as the target treatment, administration schedule, duration, and dosage.
@@
 Each derived arm includes fields such as treatment name, description, and group type, allowing for consistent representation and integration across both registry-based and publication-based sources.
-A.9 Biomarkers table
-Supplementary Figure 3: The distribution of the most frequent biomarkers under each of the five therapeutic biomarker classes.
+1.8 Biomarkers table
 Biomarkers, particularly therapeutic biomarkers (ThMARs), are critical to the design and execution of clinical trials.
@@
 Five major types of ThMARs have been identified, each playing a distinct role in drug discovery and development: (1) Pharmacodynamic Biomarkers, which ensure the pharmacological effect of a therapy; (2) Safety Biomarkers, used to assess the likelihood or severity of therapy-induced toxicity; (3) Monitoring Biomarkers, which guide ongoing clinical management by tracking patient status over time; (4) Predictive Biomarkers, which help tailor therapies to individuals most likely to benefit; and (5) Surrogate Endpoints, which serve as proxies to accelerate therapy approval.
-We leveraged TheMarker [34] as a curated reference source for therapeutic biomarkers (ThMARs) relevant to clinical trials.
+We leveraged TheMarker [10.1093/nar/gkad862] as a curated reference source for therapeutic biomarkers (ThMARs) relevant to clinical trials.
 Our processing pipeline to identify and normalize biomarker mentions consisted of two stages.
@@
 Overall, this distribution highlights the heterogeneity in biomarker utility across different therapeutic contexts, with some biomarkers, such as C reactive protein and Blood pressure, being widely applicable across multiple trial types.
-A.10 Endpoints table
-Core Outcome Measurement Sets [35, 36] provide a standardized taxonomy for classifying clinical trial outcomes, comprising five high-level domains and their associated subdomains.
+1.9 Endpoints table
+Core Outcome Measurement Sets [dodd2018taxonomy, kirkham2022core] provide a standardized taxonomy for classifying clinical trial outcomes, comprising five high-level domains and their associated subdomains.
 We adopted this COMET taxonomy to categorize outcome measurements reported in clinical trials.
@@
 The result is a structured, interpretable dataset of standardized clinical trial endpoints that facilitates cross-trial comparisons, systematic reviews, and meta-analyses.
-A.11 Adverse events table
+1.10 Adverse events table
 The Adverse Event Table is systematically populated by processing the adverseEventsModule found within the resultsSection of structured JSON records from ClinicalTrials.gov.
@@
 This systematically constructed Adverse Event Table, linked by NCTID and trial arm descriptions, serves as a valuable resource for safety analyses and complements the efficacy-focused Trial Result Table within the TrialPanorama database.
-Appendix B Benchmark building: Systematic review tasks
-B.1 Study search
+1.11 Data quality assessment
+In the database building process, a series of automated tools are used.
+Including the ones based on pattern matching with medical ontologies, the ones based on medical named entity recognition (NER) and relation extraction (RE) tools, and the ones based on large language models.
+As a result, they inevitably have introduced some errors to the built databases.
+During the data quality assessment process, we identified two common errors.
+The first type of error stems from ontology-based mapping, particularly when using tools like the RxNorm API.
+For instance, in trial NCT06425341, the drug intervention “HRS-5635 Injection” was incorrectly mapped to ‘HRS-1167” due to name similarity.
+Such mismatches are common with pattern-matching-based medical ontologies and highlight the need for careful handling when using resources like UMLS or RxNorm.
+Another common error arises from LLM-based extraction in complex fields like biomarkers.
+For example, in the clinical trial (PubMed ID: 17148749) titled “Effects of PYY1-36 and PYY3-36 on appetite, energy intake, energy expenditure, glucose and fat metabolism in obese and lean subjects”, the model incorrectly identifies “PYY” as a biomarker (e.g., PYY mRNA).
+However, a closer reading shows that “PYY” refers to the intervention being tested, not a biomarker.
+Such cases highlight the difficulty of reliably extracting biomedical concepts using LLMs.
+2 Supplementary Figures
+Supplementary Figure 1: The schematic overview of the built database.
+Supplementary Figure 2: (a) Trial outcome extraction workflow.
+Completed trials are labeled using outcome measures and matched PubMed abstracts.
+Terminated trials are categorized by mapping stop reasons to predefined outcome types via keyword similarity.
+(b) Distribution of extracted outcome types on trials from ClinicalTrials.gov.
+Supplementary Figure 3: Pipeline for drug name normalization and standardization.
+Drug names are cleaned, mapped to RxNorm, and subsequently resolved to DrugBank identifiers using a combination of indirect and direct strategies.
+Supplementary Figure 4: The distribution of the most frequent biomarkers under each of the five therapeutic biomarker classes.
+3 Supplementary Tables
+Table 1: List of data sources and the corresponding last crawl dates.
+Source Data URL Last Crawled
+ANZCTR https://www.anzctr.org.au/ April 1, 2025
+Brazilian Registry https://ensaiosclinicos.gov.br April 1, 2025
+ChiCTR https://www.chictr.org.cn/searchprojEN.html April 1, 2025
+Dutch Registry https://www.onderzoekmetmensen.nl/en/ April 1, 2025
+EU Clinical Trials https://www.clinicaltrialsregister.eu April 1, 2025
+DRKS https://drks.de/search/en/results April 1, 2025
+Iranian Registry https://irct.behdasht.gov.ir/ April 1, 2025
+ISRCTN https://www.isrctn.com April 1, 2025
+Japan Registry https://center6.umin.ac.jp/cgi-open-bin/ctr_e/ctr_view.cgi April 1, 2025
+South Korea CRIS https://cris.nih.go.kr/cris/index/index.do April 8, 2025
+Pan African Registry https://pactr.samrc.ac.za/Search_v2.aspx April 8, 2025
+Sri Lanka CTR https://slctr.lk/trials April 8, 2025
+Thai Registry https://www.thaiclinicaltrials.org April 8, 2025
+REec https://reec.aemps.es/reec/public/web.html April 9, 2025
+CTRI https://ctri.nic.in/Clinicaltrials/advancesearchmain.php April 9, 2025
+PubMed https://ftp.ncbi.nlm.nih.gov/ April 9, 2025
+Table 2: Drug mapping statistics across data sources.
+Source Total Drugs RxNorm Mapped RxNorm Rate (%) DrugBank Mapped DrugBank Rate (%)
+PubMed 953,361 579,035 60.74 241,879 25.37
+ClinicalTrials.gov 366,214 290,371 79.29 273,509 74.69
+Table 3: Statistics of the relation types in the relations table.
+Relation Type Count Percentage (%)
+PubMed - cite - PubMed 349,680 37.177
+PubMed - exclude - PubMed 237,076 25.205
+PubMed - include - PubMed 180,827 19.225
+PubMed - linked_to - ClinicalTrials.gov 167,121 17.768
+ClinicalTrials.gov - linked_to - JAPAN 2,787 0.296
+PubMed - include - ClinicalTrials.gov 954 0.101
+ClinicalTrials.gov - linked_to - KOREA 632 0.067
+PubMed - cite - ClinicalTrials.gov 425 0.045
+PubMed - exclude - ClinicalTrials.gov 406 0.043
+ClinicalTrials.gov - linked_to - GERMAN 331 0.035
+ClinicalTrials.gov - linked_to - ANZCTR 201 0.021
+ClinicalTrials.gov - linked_to - PAN_AFRICAN 56 0.006
+ClinicalTrials.gov - linked_to - SRI_LANKA 37 0.004
+ClinicalTrials.gov - linked_to - CHICTR 31 0.003
+ClinicalTrials.gov - linked_to - DUTCH 18 0.002
 Table 4: Example instance of Study Search
@@
 Ground-truth studies (PubMed ID) "20696552", "2612495", "8937535", "10563619", "2498073", "2127016", "8232944", "8112232", "17938371", "18077797", "8505632", "8453943"
-Definition & Motivation.
-This task simulates the study search step in conducting systematic reviews, where the model must comprehend a given research question and synthesize a corresponding search query to retrieve relevant studies for inclusion.
-The key motivation behind this task is to streamline a traditionally labor-intensive process that typically requires the expertise of trained librarians, particularly due to the complexity and variability of search query formats across literature databases.
-Automating this step can significantly accelerate systematic review workflows.
-An example instance is presented in Supplementary Table 7.
-Input & Output.
-The input for this task includes the setup of a systematic review, consisting of the research background, objective, and study selection criteria.
-The model is expected to generate a structured search query, which is then submitted to the PubMed Search API to retrieve the top 500 results and their associated PubMed IDs.
-A successful output should go beyond simple keyword extraction, incorporating semantic understanding and augmentation of the input to produce a comprehensive and effective query.
-Dataset Construction.
-We constructed the dataset using a relation table that links systematic reviews to primary studies.
-Specifically, we focused on instances with the relation type include, indicating that the study was selected for inclusion in the review.
-For each review, we parsed its full text to extract relevant sections and used the linked PubMed abstracts as ground-truth targets.
-This process resulted in a dataset of 8,104 examples derived from systematic review articles, covering more than 180,000 included PubMed studies.
-We reserved the most recent 100 reviews as a held-out test set.
-Evaluation Metrics.
-We evaluate this task as a retrieval problem by measuring how many ground-truth studies appear in the top search results produced by the model-generated query.
-Specifically, we report Recall@100 and Recall@500, reflecting the number of included studies retrieved among the top 100 and 500 PubMed results, respectively.
 Table 5: Example instance of Study Screening
@@
 PMID: 35708380 — [Title and abstract truncated for brevity.] PMID: 16927868 — [Title and abstract truncated for brevity.] PMID: 24464258 — [Title and abstract truncated for brevity.] — (additional candidate studies omitted for space) Screening Labels: 32382859 — Included 35708380 — Included 16927868 — Excluded
-B.2 Study screening
-Definition & Motivation.
-This task simulates the study selection step in a systematic review pipeline, where the model must assess the eligibility of candidate studies for inclusion based on the review’s predefined criteria.
-Given a systematic review’s background, objective, and eligibility criteria, along with a set of candidate studies retrieved from PubMed (each with a title and structured abstract), the model determines whether each study should be included or excluded.
-This step is often labor-intensive and requires detailed reasoning across clinical nuances and protocol designs.
-An example instance is shown in Supplementary Table 8.
-Input & Output.
-The input to this task consists of the systematic review protocol (comprising the review’s background, objective, and selection criteria) and a list of candidate studies, each represented by its title and structured abstract.
-The output is a binary label (included or excluded) for each candidate study, indicating whether it meets the review’s eligibility criteria.
-Accurate performance requires the model to match eligibility conditions against sometimes subtle or incomplete cues in the candidate study abstracts.
-Dataset Construction.
-We construct the dataset using reviews that have explicit inclusion and exclusion annotations for candidate studies, available from the relations table in our clinical trial knowledge base.
-For each review, we extract its background, objective, and selection criteria from the abstract, and retrieve associated candidate studies along with their labels (included or excluded) based on the relation types include and exclude.
-To ensure a balanced and informative evaluation, we downsample the candidate pool for each review to include a random mix of included and excluded studies, preserving diversity while controlling for class imbalance.
-Each instance contains up to 10 candidate studies, each with a title and structured abstract, paired with a single review protocol.
-Evaluation Metrics.
-We evaluate model performance using three metrics: accuracy, precision, and recall.
-Accuracy reflects the overall proportion of correctly classified studies.
-Precision measures the proportion of correctly identified included studies among all studies predicted as included, while recall measures the proportion of truly included studies that were correctly identified by the model.
-These metrics jointly assess both correctness and completeness of model screening decisions.
-B.3 Evidence summarization
-Definition & Motivation.
-This task simulates the final synthesis step of a systematic review, in which the model must draw a conclusion from the included studies’ findings.
-Given a self-contained multiple-choice question based on a Cochrane review’s objectives, the model selects the answer option that aligns with the review’s main conclusion.
-An example instance is shown in Supplementary Table 9.
-Input & Output.
-The input to the evidence summarization task consists of a self-contained multiple-choice question generated from a Cochrane review abstract (including its answer options A, B, C, …) together with the titles and full structured abstracts of all studies included in that review (ranging from 1 to 20 PMIDs).
-The model’s output is a single label (A, B, C, …) corresponding to the correct conclusion.
-Successfully performing this task requires the model to comprehend scientific text, aggregate evidence across multiple documents, and map study findings to the correct conclusion.
-Dataset Construction.
-First, Cochrane review abstracts are processed by GPT-4o using a prompt that instructs the model to generates a self-contained multiple-choice question based on the review objectives, to produce answer options covering plausible outcomes, and to output only valid JSON with the keys question, options, and answer.
-The prompt also forbids referencing the original review by phrases such as “based on” to ensure questions are fully self-contained, and the model is run at temperature 0 to guarantee deterministic outputs.
-Next, for each generated question, we retrieve the included-study PMIDs from the review metadata and load the corresponding titles and structured abstracts from our PubMed cache, discarding any entries with zero or more than twenty included studies.
-Finally, the remaining $N$ examples are sorted by numeric PMID (serving as a proxy for publication date); the most recent 1 000 examples form the test set, the preceding 500 form the validation set, and the remainder form the training set.
-Evaluation Metrics.
-We report accuracy and macro-F1 to account for class imbalance across answer choices.
 Table 6: Example instance of Evidence Summarization
@@
 Answer: C
-Appendix C Benchmark building: Trial design tasks
-C.1 Arm design
-Definition & Motivation.
-This task assesses whether the model can identify the correct treatment arms and intervention details for a given clinical trial.
-Based on the trial’s title and brief summary, the model must choose among four candidate descriptions the one that accurately represents the trial’s arm labels, arm types, and assigned interventions.
-An example instance is shown in Supplementary Table 10.
-Input & Output.
-The model is provided the official trial title and summary, followed by four arm descriptions that list the group labels, intervention types, and associated treatments.
-It must return the label corresponding to the correct description.
-Dataset Construction.
-We start with sets of up to four trial identifiers drawn from trials cited together in the same Cochrane review.
-For each identifier, a script extracts the arm groups and individual intervention details from our trial metadata.
-If fewer than four candidates are available, additional trials are added at random (with a fixed seed for reproducibility).
-To create challenging distractors, incorrect descriptions are drawn from the same arm and intervention information of other trials cited by that review.
-All items are then sorted by the numeric portion of the trial identifier and partitioned so that the latest 1,000 trials form the test set, the next 500 form validation, and the rest form training.
-Evaluation Metrics.
-We report accuracy and macro-F1 to account for class imbalance across answer choices.
+Table 7: Sample Size Estimation Question
+[…]
 Table 8: Example instance of Trial Arm Selection
 Source NCTID: NCT00000150 Question: Please select the arm or intervention descriptions that belong to the clinical trial below.
-Title: Submacular Surgery Trials (SST) Brief Summary: To determine whether surgical removal of subfoveal choroidal neovascularization (CNV) and associated hemorrhage in patients with age-related macular degeneration (AMD), the ocular histoplasmosis syndrome (OHS), or idiopathic CNV stabilizes or improves vision more often than observation.
-To determine how surgical removal compared to observation of subfoveal CNV due to AMD, OHS, or idiopathic causes changes the patient’s perception of health- and vision-related “quality of life,” as measured by telephone interview using the Medical Outcomes Survey Short Form-36 (MOS SF-36), the Hospital Anxiety and Depression Scale, and the National Eye Institute Visual Function Questionnaire (NEI VFQ-25).
+Title: Submacular Surgery Trials (SST) Brief Summary: To determine whether surgical removal of subfoveal choroidal neovascularization (CNV) and associated hemorrhage in patients with age‐related macular degeneration (AMD), the ocular histoplasmosis syndrome (OHS), or idiopathic CNV stabilizes or improves vision more often than observation.
+To determine how surgical removal compared to observation of subfoveal CNV due to AMD, OHS, or idiopathic causes changes the patient’s perception of health‐ and vision‐related “quality of life,” as measured by telephone interview using the Medical Outcomes Survey Short Form-36 (MOS SF-36), the Hospital Anxiety and Depression Scale, and the National Eye Institute Visual Function Questionnaire (NEI VFQ-25).
 To determine whether randomized trials of surgery are warranted for patients with subfoveal CNV associated with AMD not suitable for laser treatment.
@@
 Answer: C
-C.2 Eligibility criteria design
-Definition & Motivation.
-This task measures the model’s ability to match a trial’s inclusion and exclusion criteria.
-Given a trial’s title and summary, the model must select, from four text options, the one that accurately lists the trial’s eligibility rules.
-An example instance is shown in Supplementary Table 11.
-Input & Output.
-The model receives the trial title and summary along with four candidate eligibility criterion sets. It must output the correct option label.
-Dataset Construction.
-For eligibility criteria, dataset construction follows the same approach as in Arm Design: we form candidate sets from trials cited by the same Cochrane review, extract each trial’s eligibility rules from our metadata, pad with random trials when needed, and draw distractors from other cited trials to ensure relevance.
-Examples are then sorted by the numeric part of the trial identifier and split into test (latest 1,000), validation (next 500), and training (remainder).
-Evaluation Metrics.
-We report Accuracy and macro-F1 to account for class imbalance across answer choices.
 Table 9: Example instance of Trial Eligibility Criteria Selection
@@
 Inclusion: IPF diagnosis; age 40–80; FVC >=50% predicted; DLCO >=35% predicted; FVC or DLCO <=90% predicted; no improvement in past year; walk 150 m in 6 min with sat >=83% on <=6 L/min O2; Exclusion: unable PFT; obstructive lung disease or airway hyper-responsiveness; expected lung transplant within 72 weeks; active infection; liver disease; cancer or other condition likely death within 2 years; diabetes; pregnancy/lactation; substance abuse; long QT syndrome; other IPF treatment; inability to take study medication; withdrawal from IPF trials; … Answer: A
-C.3 Endpoint design
-Definition & Motivation.
-This task evaluates whether the model can identify the primary outcome measures of a clinical trial.
-Based on the trial title and summary, the model must pick the option that lists the correct primary endpoints.
-An example instance is shown in Supplementary Table 12.
-Input & Output.
-The input consists of the trial title and summary plus four outcome descriptions that concatenate the primary measures and any associated timeframes. The model must choose the correct label.
-Dataset Construction.
-For primary endpoints, dataset construction also follows the Arm Design template: trials cited together provide candidate sets, we extract each trial’s primary outcomes from our metadata, pad with random examples if needed, and sample distractor options from other cited trials.
-The resulting items are ordered by the numeric portion of the trial identifier and divided into test (latest 1,000), validation (next 500), and training splits.
-Evaluation Metrics.
-We report Accuracy and macro-F1 to account for class imbalance across answer choices.
 Table 10: Example instance of Trial Outcome Measure Selection
@@
 Visual acuity in amblyopic eye at 6 months; … Answer: D
-C.4 Sample size estimation
-Definition & Motivation.
-This task evaluates the model’s quantitative reasoning by asking it to estimate the number of participants required for a clinical trial based on its design and stated statistical assumptions.
-Success requires understanding trial objectives, translating textual assumptions (effect size, significance level, power, dropout) into a numerical enrollment target, and applying standard sample-size formulas.
-An example instance is shown in Supplementary Table 13.
-Input & Output.
-As input, the model receives: a narrative description of the trial design—official title, conditions, arm groups, interventions and primary outcomes, as well as a concise summary of the statistical assumptions extracted from the protocol text (treatment effect, alpha, power, expected dropout).
-The model must output a single integer corresponding to the estimated sample size.
-Dataset Construction & Splits.
-We linked 5,737 PubMed protocols to their NCTIDs and identified 4,315 full-text sections titled “sample size” together with 5,678 CT.gov records containing enrollment counts.
-Matching the reported CT.gov enrollment with the corresponding paper section yielded 2 143 examples where the protocol paragraph explicitly mentions the official enrollment number.
-For each example, we used a deterministic prompt to summarize the paragraph into (1) effect-size assumptions, (2) statistical parameters (significance level, power, dropout rate), and (3) the true enrollment.
-We then generated a question by combining trial metadata with these summarized assumptions.
-Finally, all examples were sorted by the numeric part of the NCTID; the most recent 1,000 form the test set, the next 500 the validation set, and the remainder the training set.
-Evaluation Metrics.
-We report (1) the mean absolute error between predicted and actual enrollments, and (2) the proportion of trials for which the model’s estimate lies within ±20 % of the true sample size.
-Table 13: Sample Size Estimation Question
-Source NCTID: NCT06330298 Question: Here is the design of a clinical trial: Official Title: Improving Social Cognition and Social Behaviour in Various Brain Disorders.
-Included conditions: Stroke; Multiple Sclerosis; Brain Tumor.
-Study Arm groups: Experimental condition: Receives T-ScEmo Treatment (type EXPERIMENTAL; receives T-ScEmo between T0 and T1; interventionNames: [“Behavioral: Treatment social cognition and emotion regulation (T-ScEmo)”]); Waiting list group: Will be on waiting list instead of treatment (type NO INTERVENTION; duration T0–T1).
-Interventions: Type BEHAVIORAL; name Treatment social cognition and emotion regulation (T-ScEmo); description T-ScEmo is a multifaceted protocol to improve social cognition, regulation of social behaviour and participation (Westerhof et al., 2017, 2019); 15×1 h live sessions + 5 online practice sessions; involves significant other and homework; armGroupLabels: [Experimental condition: Receives T-ScEmo Treatment].
-Primary Outcomes: Change in social behaviour examined by proxy; difference T0→T2 on the Dysexecutive Questionnaire Social scales proxy version (DEX-Socproxy; Spikman et al., 2013; Westerhof-Evers, 2017); time frame: through completion (8–10 months).
-Please estimate the sample size based on: significance level ($\alpha=0.05$), power ($0.801-\beta=$ 0.80), and the group×$\times$×time interaction in a repeated measures analysis.
-Answer: 84 Explanation: The study of Westerhof-Evers et al.
-[9] found that the DEX-Socproxy had an effect size of Cohen’s d = 0.2 (Cohen’s f = 0.10) for the comparison between T0 and T2.
-From these data, a small to medium effect size of Cohen’s f = 0.15 was estimated for the current study.
-Therefore, the sample size is based on a power calculation for the main outcome measure (DEX-Socproxy) regarding the condition $\times$ time interaction (T0, T2) using repeated measures analysis with effect size Cohen’s f = 0.15, $\alpha=0.05$, and power = 0.80 [29], resulting in a sample size of $n=62$.
-However, to ensure balanced distribution across patient groups, a total sample size of $n=84$ will be used (42 patients per condition).
-C.5 Trial completion assessment
-Definition & Motivation.
-This task simulates a real-world feasibility analysis during clinical trial planning, where the model must assess whether a trial is likely to reach completion or be terminated early based on its protocol design.
-In addition, if a trial is predicted to terminate, the model is asked to identify the most likely reason for termination from a predefined set of categories (e.g., enrollment issues, safety concerns, lack of efficacy, operational problems).
-This reflects practical needs in the early design stage when sponsors assess trial viability.
-An example instance is shown in Supplementary Table 14.
-Input & Output.
-Each example consists of structured metadata and protocol text for a registered interventional study, including trial phase, number of arms, masking, intervention model, primary purpose, arm descriptions, and detailed eligibility criteria.
-The output is a binary label indicating whether the trial was completed or terminated.
-For trials labeled as terminated, the model must also predict the most likely termination reason from one of several standardized categories.
-Dataset Construction.
-We construct the dataset using completed and terminated interventional trials from ClinicalTrials.gov with known outcomes.
-Structured fields are extracted from the studies, disposition, eligibility_criteria, and arms tables in our trial database.
-Only trials with definitive outcome labels (i.e., Completed or Terminated) and non-missing eligibility criteria are retained.
-Termination reasons are normalized into broad categories using a curated taxonomy (e.g., enrollment issues, lack of efficacy, safety, sponsor decision).
-To avoid label leakage, we exclude any outcomes or termination reasons from the input text fields.
-Evaluation Metrics.
-We evaluate trial completion prediction using accuracy. For terminated trials, termination reason classification is evaluated using macro-F1 to balance across categories with different frequencies.
 Table 11: Example instance of Trial Completion Assessment